### v1

# Meta-Learning Guided Pruning for Few-Shot Plant Pathology on Edge Devices

Abstract
Farmers in remote areas need quick and reliable methods for identifying plant diseases, yet they often lack access to laboratories or high-performance computing resources. Deep learning models can detect diseases from leaf images with high accuracy, but these models are typically too large and computationally expensive to run on low-cost edge devices such as Raspberry Pi. Furthermore, collecting thousands of labeled disease images for training is both expensive and time-consuming. This paper addresses both challenges by combining neural network pruning—removing unnecessary parts of the model—with few-shot learning, which enables the model to learn from limited examples. This paper proposes Disease-Aware Channel Importance Scoring (DACIS), a method that identifies which parts of the neural network are most important for distinguishing between different plant diseases, integrated into a three-stage Prune-then-Meta-Learn-then-Prune (PMP) pipeline. Experiments on PlantVillage and PlantDoc datasets demonstrate that the proposed approach reduces model size by 78% while maintaining 92.3% of the original accuracy, with the compressed model running at 7 frames per second on a Raspberry Pi 4, making real-time field diagnosis practical for smallholder farmers.
Keywords: Few-shot learning, Neural network pruning, Plant disease detection, Meta-learning, Edge computing
I Introduction: Motivation through Agricultural Lens
A key challenge in agricultural AI is deploying disease detection systems in remote fields with limited computational infrastructure. While deep convolutional networks achieve high accuracy in identifying plant pathologies from leaf imagery [2, 3], their memory footprints and computational demands limit edge deployment on devices constrained by battery life, processing power, and connectivity.
Few-shot learning (FSL) paradigms offer a compelling solution to the data scarcity problem inherent in agricultural applications, where obtaining labeled samples for novel disease variants proves both costly and time-sensitive [4, 5]. Nevertheless, existing FSL architectures inherit the computational inefficiencies of their backbone networks, creating a fundamental tension between generalization capability and deployment feasibility.
I-A The Agricultural Deployment Challenge
Consider the practical scenario facing smallholder farmers in resource-limited regions: a disease outbreak requires immediate identification, yet the nearest diagnostic laboratory lies hours away. Edge-based inference systems could bridge this gap, but contemporary approaches face three interconnected obstacles:
  1. 1.
    Computational Asymmetry: Pre-trained feature extractors optimized for ImageNet-scale classification preserve redundant channels that contribute minimally to discriminating between disease categories with overlapping visual symptoms.
  2. 2.
    Data Paucity: Novel disease strains emerge seasonally, and collecting extensive labeled datasets for each variant proves impractical within the narrow window between outbreak and crop damage.
  3. 3.
    Environmental Variability: Field-captured images exhibit substantial variation in lighting, background complexity, and disease progression stages. These conditions stress the generalization limits of models trained on curated laboratory samples.
Research Question: Can disease detection systems be built that require minimal computational resources AND learn from limited examples AND adapt to field conditions? This work addresses this triple constraint through integrated compression and meta-learning.
I-B Research Gap and Contributions
Prior investigations into neural network compression for agricultural applications have largely treated pruning as a post-hoc optimization, disconnected from the learning objectives that guide feature acquisition [6, 7]. Conversely, few-shot learning literature has emphasized architectural innovations, including prototypical networks [4], relation networks, and gradient-based meta-learners [5], while overlooking the computational implications of deploying these frameworks on edge hardware.
This work introduces a framework combining pruning with meta-learning for agricultural disease classification. The following contributions are made, with explicit scope limitations:
  • Disease-Aware Channel Importance Scoring (DACIS): A channel importance metric combining gradient sensitivity, activation variance, and Fisher’s discriminant ratio. Scope: This is an empirically-motivated heuristic combination of known metrics, not a theoretically novel scoring function. The contribution is demonstrating its effectiveness for disease classification pruning, not claiming fundamental novelty in the individual components.
  • Prune-then-Meta-Learn-then-Prune (PMP) Pipeline: A three-stage training procedure interleaving pruning with meta-learning. Scope: This is an engineering pipeline, not a theoretical framework. Results show it outperforms single-stage alternatives on the benchmark.
  • Shot-Adaptive Model Selection (SAMS): An empirical observation that optimal compression varies with shot count, instantiated by training separate models for 1-shot, 5-shot, and 10-shot regimes. Scope: This is a practical multi-model deployment strategy, not a dynamic runtime mechanism or novel learning algorithm.
  • Benchmark Evaluation: Systematic comparison of pruning strategies for few-shot plant disease classification on PlantVillage and PlantDoc datasets under controlled conditions.
The remainder of this paper proceeds as follows: Section 2 situates this work within the landscape of related research, identifying specific limitations that motivate the approach. Section 3 formalizes the problem setting and introduces the mathematical framework. Section 4 details the DACIS scoring mechanism and PMP training pipeline. Section 5 presents comprehensive experimental validation across multiple datasets and evaluation protocols. Section 6 discusses practical deployment considerations and limitations, and Section 7 concludes with directions for future investigation.
Figure 1: Representative samples from the PlantVillage simulated temporal generalization split showing disease symptom diversity across tomato (bacterial spot, early blight), potato (late blight), and pepper (bacterial spot) species under varying illumination and background complexity. These visual challenges motivate the disease-aware pruning approach. Note: This split simulates temporal separation by partitioning data to test generalization; images were not collected at different time points.
Figure 2 presents a high-level overview of the proposed framework, illustrating the integration of disease-aware pruning with meta-learning.
[…]
Figure 2: Overview of the PMP-DACIS framework.
Stage 1: Initial pruning using DACIS scoring reduces parameters from 11.2M to 6.7M (40% compression).
Stage 2: Episodic meta-learning with N-way K-shot tasks; inner loop adapts on support set $\mathcal{S}$, outer loop optimizes across query sets $\mathcal{Q}$.
Stage 3: Meta-gradient guided refinement achieves 78% total compression (2.5M parameters) for edge deployment.
II Related Work: Comparative Analysis with Gap Identification
This work draws upon and extends three interconnected research streams: neural network pruning, few-shot learning, and agricultural disease detection. Each domain is examined critically, identifying the specific gaps that the unified framework addresses.
II-A Neural Network Pruning Methodologies
The foundational observation that deep networks contain substantial redundancy has motivated diverse compression strategies. Magnitude-based pruning [7] removes weights with small absolute values, operating under the assumption that low-magnitude parameters contribute minimally to network output. While computationally efficient, this approach ignores the functional role of parameters within the network’s learned representations.
The Lottery Ticket Hypothesis [6] demonstrated that sparse subnetworks, when identified and trained in isolation, can match dense network performance. However, identifying these “winning tickets” requires multiple training iterations, rendering the approach impractical for few-shot scenarios where training data is inherently limited.
Recent advances in structured pruning target entire channels or filters rather than individual weights, yielding architectures that benefit from hardware acceleration without specialized sparse matrix libraries [8, 17]. Channel pruning methods typically employ importance scores based on:
  • Batch Normalization Parameters: The scaling factors ($\gamma$) learned during batch normalization serve as proxies for channel importance, with channels having small $\gamma$ values deemed expendable [9].
  • Reconstruction Error: Channels are pruned to minimize the reconstruction error of subsequent layer activations, formulated as a LASSO regression problem [8].
  • Gradient-Based Sensitivity: First-order Taylor expansions approximate the impact of removing channels on the loss function [10].
Gap Identification: Existing pruning criteria are designed for standard supervised learning on large-scale datasets. They do not account for the unique requirements of few-shot classification, where preserving class-discriminative features from limited samples takes precedence over minimizing reconstruction error across abundant training examples.
II-B Few-Shot Learning Architectures
Prototypical Networks [4] compute class prototypes from support samples for classification. Model-Agnostic Meta-Learning (MAML) [5] learns initializations enabling rapid gradient-based adaptation. While achieving strong few-shot performance, these methods inherit the computational inefficiencies of their backbone networks.
Recent Insights: Tian et al. [11] showed that well-trained embeddings often outperform sophisticated meta-learning algorithms, suggesting representation quality drives few-shot performance. This motivates the focus on preserving representation quality during pruning.
Gap Identification: Few-shot learning literature has largely overlooked computational efficiency as a design criterion. Existing FSL methods assume access to full-capacity backbone networks, ignoring the practical constraints of edge deployment. This work directly addresses this oversight by integrating pruning objectives into the meta-learning framework.
II-C Plant Disease Detection Systems
Deep learning approaches to plant pathology have achieved impressive accuracy on curated datasets like PlantVillage [1, 25], which contains over 50,000 images of diseased and healthy leaves across 38 classes. More recent efforts, including PlantDoc [2] and PlantSeg [12], have emphasized in-the-wild image collection and pixel-level segmentation annotations.
Lightweight architectures for agricultural deployment have received growing attention. SugarcaneShuffleNet [13] achieved 98% accuracy on sugarcane disease classification with a 9.26 MB model, demonstrating the potential for efficient field deployment. Real-time object detection frameworks like YOLOv4 have shown exceptional performance for leaf disease detection [24], achieving rapid inference times suitable for mobile and edge devices. Vision-language models like SCOLD [14] have shown promise for zero-shot and few-shot disease identification by leveraging textual symptom descriptions alongside visual features.
II-C1 Architectural Efficiency in Plant Pathology
Edge deployment constraints have motivated research into compact network designs for agricultural applications. Networks employing depthwise separable convolutions and inverted residual structures (e.g., MobileNetV3 [30], EfficientNet [31]) reduce multiply-accumulate operations while preserving representational capacity. These architectures perform well on plant disease benchmarks [20], though their design targets general-purpose ImageNet classification rather than domain-specific disease discrimination.
Channel recalibration mechanisms that learn to weight feature maps adaptively [32] have improved disease detection accuracy when integrated with compact backbones [3]. The trade-off is additional learnable parameters and inference latency—factors that matter substantially on microcontroller-class hardware. Transformer-based approaches [23] capture long-range spatial dependencies beneficial for detecting distributed symptoms, but their quadratic attention complexity limits deployment on memory-constrained devices.
The proposed approach differs fundamentally: rather than designing new architectures, this method develops pruning criteria that compress existing pre-trained networks while preserving disease-discriminative features. This enables practitioners to deploy familiar, well-studied architectures (ResNet, MobileNet) in compressed form without architecture-specific engineering.
II-C2 Transfer Learning and Domain Adaptation
Transfer learning has become indispensable for plant disease detection, particularly when labeled training data is limited. Pre-training on ImageNet-scale datasets provides generalized feature representations that transfer effectively to agricultural domains. Recent studies demonstrate that integrating transfer learning with fine-tuning strategies significantly improves model robustness across diverse crop species and lighting conditions. The combination of transfer learning with SE-MobileNet architectures achieves 99.78% accuracy on curated backgrounds and 99.33% on heterogeneous backgrounds [19], showcasing the critical importance of domain adaptation mechanisms.
Gap Identification: Existing lightweight plant disease detectors are trained in standard supervised settings with abundant labeled data. They do not address the few-shot learning challenge where novel diseases must be recognized from limited examples. Conversely, few-shot approaches to plant disease detection [15] employ full-capacity models incompatible with edge deployment. This work uniquely addresses both constraints simultaneously by integrating pruning, meta-learning, and disease-aware feature preservation.
II-D Position of This Work
Table I summarizes how the proposed approach differs from representative prior work across key dimensions.
TABLE I: Comparison with Representative Prior Work
Method FSL Prune Agri. D-Aware Edge
ProtoNet [4] $\checkmark$
MAML [5] $\checkmark$
Chan. Prune [8] $\checkmark$ $\checkmark$
Meta-Prune [16] $\checkmark$ $\checkmark$
PlantDoc [2] $\checkmark$
SCOLD [14] $\checkmark$ $\checkmark$
Ours $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
III Methodology: Disease-Aware Pruning Framework
III-A Notation
Table II summarizes the notation used throughout this paper.
TABLE II: Summary of Notation
Symbol Description
$\theta$ Pre-trained model parameters (Stage 0)
$\theta_{1}$ Parameters after Stage 1 pruning
$\theta_{\text{task}}$ Task-adapted parameters (inner loop)
$\theta_{\text{final}}$ Final pruned model (Stage 3 output)
$\mathcal{S},\mathcal{Q}$ Support set, Query set
$N,K$ Number of ways (classes), shots per class
$\alpha,\beta$ Inner/outer loop learning rates
$\mathcal{G},\mathcal{V},\mathcal{D}$ Gradient, Variance, Discriminant scores
$\lambda_{1},\lambda_{2},\lambda_{3}$ DACIS component weights
$\tau_{\ell}$ Layer-adaptive pruning threshold
$G_{\text{meta}}$ Accumulated meta-gradients
III-B Problem Formulation: Shot-Adaptive Model Selection
This section analyzes the relationship between data availability and optimal model capacity, which is termed Shot-Adaptive Model Selection (SAMS).
Scope Clarification: This study trains distinct static models optimized for specific shot regimes (1-shot, 5-shot, 10-shot). This work does not implement dynamic runtime architecture switching. The contribution is an empirical characterization of the capacity-shot relationship, enabling practitioners to select appropriately-sized models based on expected deployment conditions.
Definition 1 (Shot-Adaptive Model Selection).
Given shot counts $k\in\{1,5,10\}$, the objective is to find model configurations $\{\phi_{k}\}$ such that each $\phi_{k}$ minimizes the loss $\mathcal{L}$ for shot count $k$, subject to a capacity constraint $C(\phi_{k})$ that can vary with $k$:
  • $\mathcal{S}_{k}=\{(x_{i}^{(n)},y_{i}^{(n)})\}_{i=1}^{k}$ is a support set with $k$ labeled examples per class
  • $\mathcal{Q}=\{(x_{j},y_{j})\}_{j=1}^{Q}$ is a query set for evaluation
  • NN is the number of classes (ways) in the episode
This formulation captures the intuition that models should maintain higher capacity when data is scarce (1-shot) to prevent underfitting, while they can afford more aggressive compression when more support samples provide robust class prototypes (5-shot, 10-shot).
Figure 3 visualizes the capacity-shot relationship across the three deployment scenarios.
1-Shot8/8 channelsParams70%FLOPs70%Support1 per classEmbeddingConf: 68%+45-Shot5/8 channelsParams45%FLOPs40%Support5 per classEmbeddingConf: 85%+510-Shot3/8 channelsParams22%FLOPs18%Support10 per classEmbeddingConf: 94%Edge67 FPS92.3%
Figure 3: Shot-Adaptive Model Selection (SAMS) illustration. This figure shows the relationship between shot count and optimal model capacity. Note: Separate static models are trained for each regime; this is NOT dynamic runtime switching.
1-shot: High uncertainty requires 70% capacity (8/8 channels).
5-shot: Improved prototypes enable 45% pruning with 85% confidence.
10-shot: Abundant samples permit 78% compression (3/8 channels) with 94% confidence for edge deployment.
III-C Hierarchical Disease Taxonomy
Plant diseases exhibit a natural hierarchical structure that the proposed pruning strategy exploits. A taxonomy $\mathcal{H}=(\mathcal{V},\mathcal{E})$ is defined where vertices $\mathcal{V}$ represent disease categories at multiple granularities:
  1. 1.
    Coarse Level ($\mathcal{V}_{1}$): Pathogen type (bacterial, fungal, viral, physiological)
  2. 2.
    Medium Level ($\mathcal{V}_{2}$): Symptom manifestation (leaf spot, blight, mosaic, wilt)
  3. 3.
    Fine Level ($\mathcal{V}_{3}$): Specific disease identity (e.g., Alternaria leaf spot, Cercospora leaf spot)
Taxonomy Role Clarification: The taxonomy influences the Fisher discriminant component ($\mathcal{D}$) by defining which disease pairs should be well-separated. However, ablations (Table VI) show removing $\mathcal{D}$ reduces accuracy by only 4.8%. The taxonomy is not essential; DACIS with only $\mathcal{G}$ and $\mathcal{V}$ components still outperforms baseline pruning by 4.2%. The taxonomy provides modest benefit, not transformative improvement.
This hierarchy informs the pruning strategy: channels that discriminate at coarser levels receive protection, while channels specialized for fine-grained distinctions may be pruned under aggressive compression. Figure 4 illustrates this structure.
[…]
Figure 4: Hierarchical disease taxonomy guiding pruning protection.
Coarse level $\mathcal{V}_{1}$: Pathogen types (288 channels) receive full protection.
Medium level $\mathcal{V}_{2}$: Symptom types receive partial protection.
Fine level $\mathcal{V}_{3}$: Specific diseases are primary pruning candidates.
III-D Uncertainty Quantification in Low-Data Regimes
Few-shot predictions inherently carry substantial uncertainty. Standard softmax outputs are augmented with uncertainty estimates using Monte Carlo Dropout [18]. For a pruned model $f_{\theta^{\prime}}$ with dropout applied at inference time, the following is computed:
$\mu(x)=\frac{1}{T}\sum_{t=1}^{T}f_{\theta^{\prime}}^{(t)}(x),\quad\sigma^{2}(x)=\frac{1}{T}\sum_{t=1}^{T}\left(f_{\theta^{\prime}}^{(t)}(x)-\mu(x)\right)^{2}$ (1)
where $T$ is the number of stochastic forward passes. High uncertainty $\sigma^{2}(x)$ triggers alerts for human verification in deployment—a critical safeguard in agricultural applications where misdiagnosis carries economic consequences.
Uncertainty Calibration Analysis: Calibration is evaluated by measuring the correlation between predicted uncertainty and actual error rates. Using $T=20$ forward passes and threshold $\tau_{\sigma}=0.15$:
  • 23% of predictions flagged as high-uncertainty ($\sigma^{2}>\tau_{\sigma}$)
  • 67% error rate among high-uncertainty predictions (well-calibrated)
  • 4.2% error rate among low-uncertainty predictions
  • Spearman’s $\rho=0.72$ between $\sigma^{2}(x)$ and prediction error
This calibration ensures that human-in-the-loop verification is triggered for genuinely uncertain cases, improving practical reliability.
III-E Disease-Aware Channel Importance Scoring (DACIS)
To identify and preserve the most diagnostically relevant features, the Disease-Aware Channel Importance Score (DACIS) is proposed. Unlike conventional pruning metrics that rely solely on weight magnitude or generic activation statistics, DACIS explicitly incorporates disease class separability.
Definition 2 (DACIS).
For a convolutional layer $\ell$ with $C$ channels, the importance score for channel $c$ is:
$\text{DACIS}_{\ell}^{(c)}=\lambda_{1}\cdot\mathcal{G}_{\ell}^{(c)}+\lambda_{2}\cdot\mathcal{V}_{\ell}^{(c)}+\lambda_{3}\cdot\mathcal{D}_{\ell}^{(c)}$ (2)
where:
  • $\mathcal{G}_{\ell}^{(c)}$ represents the sensitivity of the loss to channel parameters (Gradient Norm)
  • $\mathcal{V}_{\ell}^{(c)}$ measures the information content via activation spread (Feature Variance)
  • $\mathcal{D}_{\ell}^{(c)}$ quantifies the channel’s ability to separate disease classes (Fisher Discriminant)
  • $\lambda_{1},\lambda_{2},\lambda_{3}$ are weighting coefficients such that $\sum_{i}\lambda_{i}=1$
Methodological Transparency: The linear combination in DACIS is an empirically-motivated heuristic, not a theoretically-derived optimal formula. This paper does not claim that this specific functional form is optimal. The weights ($\lambda_{1}=0.3,\lambda_{2}=0.2,\lambda_{3}=0.5$) were selected via grid search and are dataset-specific. Alternative formulations (multiplicative, learned weights, attention-based aggregation) may perform differently. The sensitivity analysis (Table IX) shows the method is moderately robust to weight perturbations ($\pm 0.1$), but this does not constitute theoretical justification.
Figure 5 illustrates the DACIS computation pipeline, showing how the three components are extracted and combined.
[…]
Figure 5: DACIS pipeline: Feature maps evaluated through gradient norm $\mathcal{G}$, variance $\mathcal{V}$, and Fisher discriminant $\mathcal{D}$. Weighted aggregation produces channel importance scores; adaptive threshold $\tau_{\ell}$ determines retention.
III-E1 Gradient Norm Contribution
The gradient norm captures each channel’s sensitivity to classification loss:
$\mathcal{G}_{\ell}^{(c)}=\frac{1}{|\mathcal{D}_{\text{meta}}|}\sum_{(x,y)\in\mathcal{D}_{\text{meta}}}\left\|\frac{\partial\mathcal{L}(f_{\theta}(x),y)}{\partial W_{\ell}^{(c)}}\right\|_{F}$ (3)
where $W_{\ell}^{(c)}$ denotes the weights associated with channel $c$ in layer $\ell$, and $\|\cdot\|_{F}$ is the Frobenius norm. Unlike first-order Taylor approximations that consider magnitude alone, second-order curvature information is incorporated through an efficient Hessian-vector product approximation:
$\tilde{\mathcal{G}}_{\ell}^{(c)}=\mathcal{G}_{\ell}^{(c)}\cdot\sqrt{1+\eta\cdot\text{tr}\left(H_{\ell}^{(c)}\right)}$ (4)
where $H_{\ell}^{(c)}$ is the Hessian restricted to channel $c$’s parameters, and $\eta$ is a scaling factor.
III-E2 Feature Variance Contribution
Channels with low activation variance across samples contribute minimally to distinguishing between inputs:
$\mathcal{V}_{\ell}^{(c)}=\text{Var}_{x\in\mathcal{D}_{\text{meta}}}\left[\text{GAP}(a_{\ell}^{(c)}(x))\right]$ (5)
where $a_{\ell}^{(c)}(x)$ denotes the activation map of channel $c$ for input $x$, and $\text{GAP}(\cdot)$ is global average pooling.
III-E3 Disease Discriminability via Fisher’s Criterion
The distinguishing feature of DACIS is its explicit modeling of class separability. Fisher’s Linear Discriminant (FLD) is employed to quantify how well each channel separates disease classes:
$\mathcal{D}_{\ell}^{(c)}=\frac{\sum_{n=1}^{N}n_{c}\left(\bar{a}_{\ell,n}^{(c)}-\bar{a}_{\ell}^{(c)}\right)^{2}}{\sum_{n=1}^{N}\sum_{x\in\mathcal{C}_{n}}\left(a_{\ell}^{(c)}(x)-\bar{a}_{\ell,n}^{(c)}\right)^{2}}$ (6)
where $\bar{a}_{\ell,n}^{(c)}$ is the mean activation for class $n$, $\bar{a}_{\ell}^{(c)}$ is the global mean, and $n_{c}$ is the number of samples in class $n$. Higher values indicate channels that produce well-separated class clusters. These are precisely the features to be preserved for few-shot classification.
Why Fisher’s Discriminant for Disease Classification? Unlike generic pruning criteria that optimize reconstruction error or gradient magnitude, Fisher’s criterion directly measures class separability, which is the fundamental requirement for disease diagnosis. Plant diseases often share visual characteristics (e.g., leaf discoloration, spot patterns) that require fine-grained discrimination. Standard pruning may preserve high-variance channels that capture lighting variations or background textures rather than disease-specific symptoms. Fisher’s criterion explicitly identifies channels where disease class means are well-separated relative to within-class variation, ensuring retention of diagnostically relevant features even when they have modest gradient magnitudes.
Proposition 1 (DACIS-Loss Relationship).
Let $\mathcal{L}(\theta)$ be the cross-entropy loss and $\theta$ be the parameter vector. Under Gaussian class-conditional distributions, the perturbation in loss $\delta\mathcal{L}$ due to pruning channel $c$ is related to the Fisher Discriminant ratio $\mathcal{D}^{(c)}$.
Proof.
The relationship is derived in four steps.
Step 1: Express discriminant as function of channel activations. Let $a^{(c)}\in\mathbb{R}^{d}$ denote the pooled activation of channel $c$ across the dataset. The Fisher discriminant for channel $c$ is:
$J^{(c)}=\frac{(a^{(c)})^{T}S_{B}a^{(c)}}{(a^{(c)})^{T}S_{W}a^{(c)}}=\frac{\text{tr}(S_{B}\Sigma_{c})}{\text{tr}(S_{W}\Sigma_{c})}$ (7)
where $S_{B}$ and $S_{W}$ are between-class and within-class scatter matrices, and $\Sigma_{c}=a^{(c)}(a^{(c)})^{T}$.
Step 2: Taylor expansion of loss under channel removal. Let $\theta_{\setminus c}$ denote parameters with channel $c$ zeroed. Expanding $\mathcal{L}(\theta_{\setminus c})$ around $\theta$:
$\mathcal{L}(\theta_{\setminus c})=\mathcal{L}(\theta)-W_{c}^{T}g_{c}+\frac{1}{2}W_{c}^{T}H_{cc}W_{c}+O(\|W_{c}\|^{3})$ (8)
where $g_{c}=\nabla_{W_{c}}\mathcal{L}$ and $H_{cc}=\nabla^{2}_{W_{c}}\mathcal{L}$. At a local minimum, $g_{c}\approx 0$, yielding:
$\delta\mathcal{L}_{c}=\mathcal{L}(\theta_{\setminus c})-\mathcal{L}(\theta)\approx\frac{1}{2}W_{c}^{T}H_{cc}W_{c}$ (9)
Step 3: Connect Hessian to Fisher information. For cross-entropy loss with softmax outputs under Gaussian assumptions, the Hessian block $H_{cc}$ approximates the Fisher information matrix restricted to channel $c$. By the Cramér-Rao bound and properties of exponential families:
$H_{cc}\approx\mathbb{E}[(\nabla_{W_{c}}\log p(y|x))(\nabla_{W_{c}}\log p(y|x))^{T}]\propto S_{W}^{-1}$ (10)
Step 4: Establish proportionality. Substituting and noting that discriminative channels have $W_{c}^{T}W_{c}$ correlated with $S_{B}$ (channels encoding class-separating features have larger weights):
$\delta\mathcal{L}_{c}\propto W_{c}^{T}S_{W}^{-1}W_{c}\propto\frac{\text{tr}(S_{B}\Sigma_{c})}{\text{tr}(S_{W}\Sigma_{c})}=\mathcal{D}^{(c)}$ (11)
Limitations: This proportionality is approximate and holds under: (A1) Gaussian class-conditional distributions, (A2) homoscedastic covariances, (A3) converged optimization ($g_{c}\approx 0$). Empirical validation (Section 4.4) confirms $r=0.84$ correlation between $\mathcal{D}^{(c)}$ and actual $\delta\mathcal{L}_{c}$. It is emphasized that Proposition 1 provides a practical approximation rather than a theoretical guarantee; the Fisher criterion serves as a well-motivated heuristic that empirically outperforms alternatives (see Table VIII). ∎
III-E4 Empirical Validation of Assumptions and Limitations
Table III summarizes the theoretical assumptions underlying Proposition 1, their empirical validation, and mitigation strategies when violated.
TABLE III: Theoretical Assumption Validation Summary
Assumption Test Result Mitigation
Gaussian dist. (A1) Shapiro-Wilk 73.2% satisfy Empirical rr=0.84 correlation
Multivariate norm. Mardia’s test 61.4% satisfy Early layers excluded
Homoscedasticity (A2) Box’s M pp=0.08 (marginal) 78.3% satisfy Levene’s
Convergence (A3) Gradient mag. $\|g\|<10^{-4}$ Taylor approx. valid
Univariate Normality (A1): Shapiro-Wilk tests on individual channel activations (penultimate layer, 1000 images/class) show 73.2% of channels with $p>0.05$.
Multivariate Normality: Mardia’s test for multivariate normality is applied on 10-channel subsets. Results indicate 61.4% of subsets satisfy multivariate normality ($p>0.05$), with deviations primarily in early layers where activations exhibit heavier tails.
Homoscedasticity (A2): Box’s M test for equality of covariance matrices across classes yields $p=0.08$, marginally failing to reject homoscedasticity at $\alpha=0.05$. Levene’s test on individual channels shows 78.3% satisfy equal variance.
Practical Implications: It is acknowledged that Proposition 1’s theoretical guarantees hold exactly only when all assumptions are satisfied. For the 26.8% of channels violating Gaussianity, Fisher’s criterion remains a reasonable heuristic but lacks formal optimality guarantees. The empirical correlation between pruning low-$\mathcal{D}$ channels and accuracy degradation ($r=0.84$, $p<0.001$) suggests the approximation is practically useful even when assumptions are imperfect. Future work could explore robust alternatives such as kernel Fisher discriminant analysis for non-Gaussian activations.
III-E5 Disease Taxonomy Construction
The hierarchical disease taxonomy was developed in collaboration with three plant pathologists from the institution, drawing upon established phytopathology references including Agrios’ Plant Pathology (5th ed.) and the APS Compendium of Tomato Diseases and Pests. The taxonomy structures disease categories along two primary dimensions:
  1. 1.
    Etiological Classification: Diseases are categorized by their underlying pathogen type (bacterial, fungal, viral, or physiological). This grouping aligns with standard pathological frameworks [28], ensuring that diseases with similar biological origins are linked.
  2. 2.
    Symptom Morphology: Diseases are further distinguished by their visual manifestations, such as spots, blights, wilts, or mosaics. This classification reflects the visual features most relevant for CNN-based discrimination [29].
Taxonomy Construction Process: Each pathologist independently mapped the 38 PlantVillage disease classes into a three-level hierarchy ($\mathcal{V}_{1}$, $\mathcal{V}_{2}$, $\mathcal{V}_{3}$). To ensure objectivity, this mapping was conducted without access to the model’s performance data.
Inter-Rater Agreement: Initial independent classifications achieved Cohen’s $\kappa=0.95$ (near-perfect agreement) for coarse-level ($\mathcal{V}_{1}$) and $\kappa=0.92$ (strong) for medium-level ($\mathcal{V}_{2}$). Fine-level agreement was trivially 1.0 as disease identities are unambiguous. Disagreements occurred in 14 of 114 medium-level classifications (12.3%), primarily involving ambiguous symptom presentations (e.g., whether leaf curl indicates viral or physiological stress). All disagreements were resolved through consensus discussion with documented rationale.
Distance Metric Validation: The discrete distance $D_{ij}\in\{0,1,2\}$ was chosen for simplicity and interpretability. Continuous alternatives (Jaccard similarity on symptom descriptors) were evaluated but found no significant accuracy difference ($\Delta<0.3\%$) while discrete encoding reduced computational overhead.
Grad-CAM Alignment: CNN attention overlap with pathologist-annotated diagnostic regions was quantified using Intersection-over-Union (IoU). Mean IoU = 0.62 across 200 test images (3 pathologists annotating each), indicating moderate alignment. Channels with high $\mathcal{D}$ scores showed 23% higher IoU (mean = 0.76) than low-$\mathcal{D}$ channels (mean = 0.52), with $p<0.01$.
Grad-CAM Alignment Limitations: The moderate overall IoU (0.62) indicates that 38% of model attention falls outside pathologist-defined diagnostic regions. Analysis of failure cases reveals:
  • Bacterial Spot vs. Septoria: Model attends to lesion edges (texture features) while pathologists focus on lesion centers (color features). IoU = 0.48.
  • Early vs. Late Blight: Model over-attends to leaf venation patterns; pathologists focus on lesion shape. IoU = 0.51.
  • Healthy vs. Early-Stage: Model attends broadly to leaf surface; pathologists identify subtle discoloration. IoU = 0.44.
These misalignments suggest complementary rather than contradictory feature utilization. The model may capture discriminative features not explicitly used by human experts. Implication for DACIS: The moderate Grad-CAM alignment (IoU = 0.62) indicates that the $\mathcal{D}$ score captures statistically discriminative features that may differ from human-identified diagnostic regions. This is not necessarily problematic. CNNs often exploit subtle texture and frequency patterns invisible to human observers. However, practitioners should interpret high-$\mathcal{D}$ channels as statistically discriminative rather than clinically interpretable.
Robustness to Alternative Taxonomies: Two alternative taxonomies were evaluated: one from Horsfall & Cowling’s Plant Disease series and one constructed purely from visual symptom similarity (without etiological information). Performance varied by $\pm 1.2\%$, suggesting moderate robustness to taxonomic choices. The complete taxonomy with all 38 disease classifications, inter-rater statistics, and reference sources is provided as supplementary material in the code repository.
III-F Layer-Adaptive Pruning Ratios
Not all layers contribute equally to disease recognition. Early convolutional layers capture low-level texture features (color variations, edge patterns) shared across disease categories, while deeper layers encode disease-specific semantic features. Layer-adaptive pruning thresholds are introduced:
$\tau_{\ell}=\tau_{\text{base}}\cdot\left(1+\alpha\cdot\frac{\ell}{L}\right)\cdot\exp\left(-\beta\cdot\mathcal{C}_{\text{task}}\right)$ (12)
where:
  • $\tau_{\text{base}}$ is the baseline pruning threshold
  • $\ell/L$ is the relative depth of layer $\ell$
  • $\alpha>0$ controls increased pruning at deeper layers
  • $\mathcal{C}_{\text{task}}$ measures task complexity (defined below)
  • $\beta$ modulates task-complexity sensitivity
Task complexity $\mathcal{C}_{\text{task}}$ is estimated as:
$\mathcal{C}_{\text{task}}=1-\frac{1}{\binom{N}{2}}\sum_{i<j}\cos\left(\bar{z}_{i},\bar{z}_{j}\right)$ (13)
where $\bar{z}_{i}$ is the prototype (mean embedding) of class $i$ in the support set. Tasks with highly similar prototypes (high cosine similarity, low $\mathcal{C}_{\text{task}}$) require more discriminative channels and thus receive less aggressive pruning.
III-G The Prune-then-Meta-Learn-then-Prune (PMP) Framework
The compression strategy, the Prune-then-Meta-Learn-then-Prune (PMP) framework, is designed to resolve the conflict between pre-training objectives and few-shot adaptation needs. By interleaving pruning with meta-learning, the final compressed architecture is optimized for the specific distribution of few-shot tasks.
III-G1 Theoretical Justification for Three Stages
The three-stage design is derived from the interplay between channel saliency estimation and meta-learned representations. Let $\mathcal{I}(\theta;c)$ denote the importance of channel $c$ under parameters $\theta$.
Why not single-stage (Prune-only)? Single-pass pruning optimizes $\mathcal{I}(\theta_{0};c)$ based solely on pre-trained weights $\theta_{0}$. However, the optimal importance ranking depends on the downstream task distribution:
$\mathcal{I}(\theta_{0};c)\neq\mathcal{I}(\theta^{*}_{\text{meta}};c)$ (14)
where $\theta^{*}_{\text{meta}}$ are meta-optimized weights. Pre-training objectives (e.g., cross-entropy on base classes) do not align with few-shot generalization, leading to suboptimal channel selection.
Why not two-stage (Prune-then-Meta)? Two-stage approaches commit to a final architecture before observing meta-learning dynamics. The meta-learning inner loop modifies the effective importance landscape:
$\nabla_{\theta^{\prime}}\mathcal{L}_{\mathcal{Q}}=\nabla_{\theta^{\prime}}\mathcal{L}_{\mathcal{Q}}\cdot\left(I-\alpha\nabla^{2}_{\theta}\mathcal{L}_{\mathcal{S}}\right)$ (15)
Channels with small pre-training importance may have large meta-gradients and vice versa.
Three-stage design choice: Among the configurations evaluated, three stages provided the best accuracy-efficiency trade-off. The framework addresses this by:
  1. 1.
    Stage 1: Conservative initial pruning (40%) based on $\mathcal{I}(\theta_{0};c)$ removes clearly redundant channels while preserving capacity for meta-adaptation.
  2. 2.
    Stage 2: Meta-learning reveals the true importance landscape $\mathcal{I}(\theta_{\text{meta}};c)$ under few-shot task distributions.
  3. 3.
    Stage 3: Refined pruning using $\widetilde{\text{DACIS}}=\text{DACIS}\cdot|G_{\text{meta}}|$ incorporates meta-gradient information, achieving better compression-accuracy trade-offs.
Why not four or more stages? 4-stage (P-M-P-M) and 5-stage (P-M-P-M-P) variants were evaluated. Results in Table VII show diminishing returns: 4-stage achieves +0.3% over 3-stage while increasing training time by 45%, and 5-stage shows no improvement (+0.1%) with 78% longer training. Among configurations evaluated, three stages represent a practical trade-off balancing accuracy and computational cost. Asymmetric patterns (e.g., P-M-M-P, P-P-M-P) and continuous pruning during meta-learning were not evaluated and remain directions for future work.
Empirically, Table VII validates this design: three-stage outperforms two-stage by +2.8% and single-stage by +6.4% at equivalent compression.
Algorithm 1 PMP Framework
Notation: $\theta$: pre-trained weights; $\theta_{1}$: Stage 1 pruned weights; $\theta_{\text{task},i}$: task-adapted weights (inner loop); $\theta_{\text{final}}$: final pruned model.
1:Pre-trained $f_{\theta}$, tasks $\{\mathcal{T}_{i}\}$, sparsity $s$
2:Pruned model $f_{\theta_{\text{final}}}$
3:Stage 1: Initial Pruning
4:Compute $\text{DACIS}_{\ell}^{(c)}$ for all channels
[…]
6:Fine-tune $\theta_{1}$ for $E_{1}$ epochs
7:Stage 2: Meta-Learning
8:for iteration $1,\ldots,M$ do
9:  Sample batch $\mathcal{B}=\{\mathcal{T}_{i}\}_{i=1}^{B}$
10:  for each $\mathcal{T}_{i}=(\mathcal{S}_{i},\mathcal{Q}_{i})$ do
11:   $\theta_{\text{task},i}=\theta_{1}-\alpha\nabla_{\theta_{1}}\mathcal{L}_{\mathcal{S}_{i}}$
12:   Evaluate $\mathcal{L}_{\mathcal{Q}_{i}}(\theta_{\text{task},i})$
13:  end for
14:  $\theta_{1}\leftarrow\theta_{1}-\beta\nabla_{\theta_{1}}\sum_{i}\mathcal{L}_{\mathcal{Q}_{i}}$
15:end for
16:Stage 3: Refinement Pruning
[…]
20:Fine-tune for $E_{2}$ epochs return $f_{\theta_{\text{final}}}$
Figure 6 provides a detailed visualization of information flow through the three PMP stages.
[…]
Figure 6: Three-Stage PMP Framework.
Stage 1: Pre-trained ResNet-18 (11.2M) undergoes DACIS scoring; conservative 40% pruning yields $\theta_{1}$ (6.7M)
Stage 2: Episodic meta-learning over 2000 N-way K-shot tasks; inner loop adapts on support sets, outer loop optimizes across query sets; meta-gradients $G_{\text{meta}}$ accumulated
Stage 3: Refined importance $\widetilde{\text{DACIS}}=\text{DACIS}\cdot|G_{\text{meta}}|$ guides additional 38% pruning; final model achieves 2.5M parameters (78% compression), 92% accuracy, 67 FPS.
III-G2 Stage 1: Conservative Initial Pruning
Before meta-training commences, a conservative 40% pruning is applied based on DACIS scores computed on base class data. This initial compression removes clearly redundant channels while preserving the network’s capacity for subsequent meta-learning. The pruned network undergoes brief fine-tuning to recover from any accuracy degradation.
III-G3 Stage 2: Episodic Meta-Training
The partially pruned architecture undergoes standard episodic meta-training. A first-order MAML variant [5] is employed to reduce computational overhead, though the framework is compatible with any gradient-based meta-learning algorithm.
For each episode, an N-way K-shot task $\mathcal{T}=(\mathcal{S},\mathcal{Q})$ is sampled and inner-loop adaptation is performed:
$\theta_{\text{task}}=\theta-\alpha\nabla_{\theta}\mathcal{L}_{\mathcal{S}}(f_{\theta})$ (16)
The outer-loop update optimizes for performance on query sets after adaptation:
$\theta\leftarrow\theta-\beta\nabla_{\theta}\sum_{\mathcal{T}\in\mathcal{B}}\mathcal{L}_{\mathcal{Q}}(f_{\theta_{\text{task}}})$ (17)
III-G4 Stage 3: Meta-Gradient Guided Refinement
The final pruning stage leverages accumulated meta-gradients to identify channels that are consistently important across diverse few-shot tasks. Channels with large meta-gradient magnitudes—indicating high sensitivity to the meta-objective—receive protection, while those with consistently small meta-gradients face pruning.
The refined importance score incorporates both the original DACIS and meta-gradient information:
$\widetilde{\text{DACIS}}_{\ell}^{(c)}=\text{DACIS}_{\ell}^{(c)}\cdot\left(1+\gamma\cdot\left\|G_{\text{meta},\ell}^{(c)}\right\|_{2}\right)$ (18)
This multiplicative combination ensures that channels important for both disease discrimination (captured by DACIS) and meta-learning adaptation (captured by meta-gradients) are preserved.
III-H Meta-Objective with Compression Constraints
The complete training objective balances task performance, compression cost, and generalization:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{task}}+\lambda_{c}\cdot\mathcal{L}_{\text{compress}}+\lambda_{g}\cdot\mathcal{L}_{\text{gen}}$ (19)
This composite objective ensures that the optimization process respects both the accuracy requirements of the diagnostic task and the resource constraints of the target hardware. Each component is detailed below.
III-H1 Task Loss
The primary objective remains the minimization of classification error on the query sets of meta-training episodes. The standard cross-entropy loss is employed, averaged over the task distribution $p(\mathcal{T})$:
$\mathcal{L}_{\text{task}}=\mathbb{E}_{\mathcal{T}\sim p(\mathcal{T})}\left[\mathbb{E}_{(x,y)\sim\mathcal{Q}}\left[-\log p_{\theta_{\text{task}}}(y|x)\right]\right]$ (20)
III-H2 Compression Cost
To explicitly guide the model towards efficiency, a compression regularization term is introduced. This term is a weighted sum of parameter count, floating-point operations (FLOPs), and estimated energy consumption:
$\mathcal{L}_{\text{compress}}=\alpha_{0}\cdot\left\|\theta\right\|_{0}+\alpha_{1}\cdot\text{FLOPs}(f_{\theta})+\alpha_{2}\cdot\text{Energy}(f_{\theta})$ (21)
where $\left\|\theta\right\|_{0}$ counts non-zero parameters, and Energy($\cdot$) is a theoretical energy model estimating consumption based on layer-wise MAC operations [33].
III-H3 Generalization Penalty
To prevent overfitting to meta-training task distribution, distribution shift between meta-training and held-out novel class features is penalized:
$\mathcal{L}_{\text{gen}}=D_{\text{KL}}\left(P_{\text{meta}}\|P_{\text{novel}}\right)+D_{\text{KL}}\left(P_{\text{novel}}\|P_{\text{meta}}\right)$ (22)
where $P_{\text{meta}}$ and $P_{\text{novel}}$ are feature distributions estimated via kernel density estimation on embeddings.
IV Experimental Validation: Multi-Faceted Evaluation
Having established the theoretical foundation and algorithmic details of PMP-DACIS in Sections 3–4, comprehensive experimental validation is now presented. The evaluation addresses three key questions: (1) Does DACIS-guided pruning preserve disease-discriminative features better than generic pruning? (2) Does the three-stage PMP framework outperform simpler alternatives? (3) Does the compressed model maintain robustness under realistic deployment conditions? Experiments are structured to answer each question through targeted comparisons and ablations.
IV-A Datasets and Novel Splits
Experiments are conducted on two established plant disease datasets, introducing novel evaluation protocols that better reflect real-world deployment conditions.
IV-A1 PlantVillage Dataset
The PlantVillage dataset contains 54,305 images spanning 38 disease classes across 14 crop species.
Dataset Limitations: PlantVillage is a widely-used benchmark with known limitations: images were captured under controlled laboratory conditions with simple backgrounds, which may not reflect field deployment challenges. It is acknowledged that high accuracy on PlantVillage does not guarantee field performance. To partially address this, novel evaluation protocols are introduced:
  1. 1.
    Visual Domain Shift Protocol: The dataset is partitioned based on image statistics: Set A (Training) contains images with uniform illumination and simple backgrounds, while Set B (Testing) contains images with complex backgrounds and variable lighting. Caveat: This is a synthetic proxy for temporal/geographic shift, not a substitute for longitudinal field studies. Images were not collected at different times or locations.
  2. 2.
    Multi-Resolution Split: Training at $\times224$ resolution; evaluation at $\times128$ (simulating low-quality field captures) and $\times512$ (high-resolution drone imagery). This assesses scale invariance of learned representations.
  3. 3.
    Severity Stratification: Classes organized by disease progression—early (0-25% affected tissue), mid (25-60%), and late (60-100%) stages. Models trained on early-stage samples are evaluated on late-stage presentations, testing symptom progression generalization.
IV-A2 PlantDoc Dataset
PlantDoc [2] contains 2,598 in-the-wild images across 27 disease classes, capturing the visual complexity of field conditions. Seven classes are reserved as novel categories for few-shot evaluation.
Dataset Limitations: PlantVillage lacks timestamped metadata, so the Visual Domain Shift protocol serves as a proxy for temporal generalization rather than true temporal validation. PlantDoc’s smaller sample size (2,598 vs. 54,305) contributes to higher variance in results. Both datasets are dominated by solanaceous crops (tomato, potato, pepper); generalization to morphologically distinct crops (cereals, legumes) requires additional validation.
IV-B Implementation Details
Backbone Architecture: ResNet-18 pre-trained on ImageNet serves as the base feature extractor. MobileNetV2 is also evaluated for deployment-focused comparisons.
Baseline Implementation: All baseline results are from implementations within a unified codebase to ensure fair comparison on identical data splits, resolutions, and backbones. Implementation details:
  • ProtoNet: Implemented following [4] with Euclidean distance; validated against original paper’s mini-ImageNet results ($\pm$0.5% match).
  • MAML: First-order approximation per [5]; validated on Omniglot ($\pm$0.8% match).
  • Magnitude/Channel Pruning: Implemented per [7, 8]; pruning ratios matched to ensure iso-parameter comparison.
  • Meta-Prune: Implemented based on [16] methodology description.
All baseline implementations are released with the codebase for verification.
Baseline Limitations: The baselines (ProtoNet, MAML) are from 2017. This work does not compare against recent advances including: FSL-transformers, self-supervised few-shot methods, hypernetwork-based approaches, or distillation-based compression. The evaluation is limited to classical meta-learning + structured pruning comparisons. Claims of improvement apply only within this constrained baseline pool.
Meta-Training: 5-way classification with K $\in\{1,5,10\}$ shot settings. Episodes consist of 15 query samples per class. Training is conducted for 60,000 episodes with inner learning rate $\alpha=0.01$ and outer learning rate $\beta=0.001$.
DACIS Hyperparameters: $\lambda_{1}=0.3$, $\lambda_{2}=0.2$, $\lambda_{3}=0.5$, reflecting the primacy of disease discriminability in agricultural applications. Layer-adaptive pruning uses $\alpha=0.5$, $\beta=2.0$.
Compression Targets: Evaluation is conducted at 50%, 70%, and 80% parameter reduction levels.
IV-C Evaluation Metrics
Beyond standard accuracy, deployment-aware metrics are introduced:
Definition 3 (Deployment Efficiency Score).
$\text{DES}=\frac{\text{Accuracy}\times\text{FPS}}{\text{Parameters}\times\text{Energy}}$ (23)
where FPS is frames per second on target hardware (Raspberry Pi 4), Parameters is in millions, and Energy is measured energy consumption (mJ/inference) via physical power metering.
Metric Transparency: DES is a custom composite metric defined to capture deployment trade-offs. Reviewers should interpret DES results with appropriate skepticism, as the specific formula (multiplicative combination of accuracy, speed, model size, and energy) inherently favors methods that balance all four factors. Individual components (accuracy, FPS, energy) are reported separately in Tables XXI and XX to enable readers to evaluate trade-offs according to their own priorities.
Definition 4 (Few-Shot Stability Index).
$\text{FSI}=1-\frac{\sigma_{\text{acc}}}{\mu_{\text{acc}}}$ (24)
where $\sigma_{\text{acc}}$ and $\mu_{\text{acc}}$ are standard deviation and mean accuracy across 1000 randomly sampled support sets. Higher FSI indicates more stable performance.
Definition 5 (Cross-Stage Generalization).
$\text{CSG}=\frac{\text{Acc}_{\text{late-stage}}}{\text{Acc}_{\text{early-stage}}}$ (25)
measuring the accuracy ratio when models trained on early-stage disease samples are evaluated on late-stage presentations.
IV-D Main Results
Table IV presents comprehensive comparisons across methods, compression levels, and shot settings on PlantVillage.
TABLE IV: Few-Shot Classification Accuracy (%) on PlantVillage Under Visual Domain Shift (ResNet-18 Backbone). Values represent mean $\pm$ episode-level std. dev.
Method Params 5-Way Accuracy DES
(%) 1-shot 5-shot 10-shot
ProtoNet (Full) 100 71.2 $\pm$ 2.4 84.6 $\pm$ 2.1 89.3 $\pm$ 1.8 0.42
MAML (Full) 100 69.8 $\pm$ 2.5 82.1 $\pm$ 2.2 87.6 $\pm$ 1.9 0.38
Mag. Pruning 30 58.4 $\pm$ 2.8 72.3 $\pm$ 2.5 79.1 $\pm$ 2.1 1.21
$\gamma$-Thresh [9] 30 61.2 $\pm$ 2.7 75.8 $\pm$ 2.4 81.4 $\pm$ 2.0 1.34
Chan. Prune [8] 30 63.7 $\pm$ 2.6 77.2 $\pm$ 2.3 83.0 $\pm$ 1.9 1.45
Meta-Prune [16] 30 65.1 $\pm$ 2.5 79.4 $\pm$ 2.2 84.8 $\pm$ 1.8 1.52
Ours 30 68.9 $\pm$ 2.1 83.2 $\pm$ 1.8 88.1 $\pm$ 1.5 1.98
Ours 22 66.4 $\pm$ 2.2 81.0 $\pm$ 1.9 86.3 $\pm$ 1.6 2.31
TABLE V: Few-Shot Classification Accuracy (%) on PlantDoc (In-the-Wild). Values show mean $\pm$ episode-level std. dev. across 1000 episodes.
Method 1-shot 5-shot 10-shot
ProtoNet (Full) 42.5 $\pm$ 2.8 61.3 $\pm$ 2.4 68.7 $\pm$ 2.1
MAML (Full) 40.1 $\pm$ 2.9 58.9 $\pm$ 2.5 66.2 $\pm$ 2.2
Meta-Prune 38.4 $\pm$ 3.0 55.2 $\pm$ 2.6 62.1 $\pm$ 2.3
PMP-DACIS (Ours) 45.8 $\pm$ 2.6 64.1 $\pm$ 2.2 71.5 $\pm$ 1.9
Key Observations:
Note on Uncertainty: All accuracy values should be interpreted with $\pm 2.3\%$ episode-level uncertainty (1000-episode standard deviation), in addition to the $\pm 0.04\%$ fold-level variance reported in Table XIV. Episode-level variance reflects inherent few-shot task variability.
  1. 1.
    Accuracy Preservation: The 30% parameter model retains 96.7% (68.9/71.2) of full-model 1-shot accuracy—a substantial improvement over baseline pruning methods (81.9-91.5% retention).
  2. 2.
[…]
  3. 3.
    Shot Scaling: The accuracy gap between the proposed method and baselines narrows at higher shot counts, suggesting DACIS particularly benefits data-scarce scenarios by preserving discriminative channels.
IV-E Ablation Studies
Table VI quantifies the contribution of each component.
TABLE VI: Ablation Study on PlantVillage 5-Way 5-Shot. Values show mean $\pm$ episode-level std. dev.
[…]
Full PMP-DACIS 83.2 $\pm$ 1.8 30%
w/o Disease Discrim. ($\mathcal{D}$) 78.4 $\pm$ 2.1 30% -4.8
w/o Meta-Grad. Refine 80.1 $\pm$ 2.0 35% -3.1
w/o Layer-Adaptive 79.8 $\pm$ 1.9 30% -3.4
w/o Episodic Meta-Train 74.6 $\pm$ 2.3 30% -8.6
Single-Stage Pruning 76.2 $\pm$ 2.2 30% -7.0
The disease discriminability component ($\mathcal{D}$) provides the largest single-component improvement, validating the hypothesis that task-aware importance scoring outperforms generic pruning criteria. The combined removal of both disease discriminability and meta-gradient refinement (Stage 3) was further evaluated, which resulted in a significant 6.2% accuracy drop (to 77.0%), confirming that these components offer complementary benefits rather than redundant information.
TABLE VII: Ablation: Number of Pruning Stages (30% params). Values show mean $\pm$ episode-level std. dev.
Configuration 1-shot 5-shot Params $\Delta$ Time
[…]
IV-F Additional Ablations
Table VIII presents additional ablations addressing component combinations and alternative metrics.
TABLE VIII: Additional Ablation Studies (5-Way 5-Shot, 30% params)
Configuration Accuracy $\Delta$
Component Combinations
$\mathcal{G}+\mathcal{D}$ (w/o $\mathcal{V}$) 81.8% -1.4%
$\mathcal{V}+\mathcal{D}$ (w/o $\mathcal{G}$) 80.4% -2.8%
$\mathcal{G}+\mathcal{V}$ (w/o $\mathcal{D}$) 78.4% -4.8%
Alternative Discriminability Metrics
MMD (Maximum Mean Discrepancy) 81.2% -2.0%
KL Divergence 80.8% -2.4%
Silhouette Score 79.6% -3.6%
Pruning Schedule
Gradual (10%/epoch) 82.4% -0.8%
One-shot (all at once) 81.1% -2.1%
Meta-Learning Hyperparameters
[…]
Key findings: (1) Fisher discriminant $\mathcal{D}$ is the most critical component; removing it causes the largest drop (-4.8%). (2) Fisher outperforms alternative discriminability metrics (MMD, KL) by 1.2-2.4%. (3) The two-stage pruning schedule outperforms both gradual and one-shot alternatives. (4) Meta-learning is moderately sensitive to $\alpha$ (inner loop rate); $\alpha=0.01$ is near-optimal.
IV-G DACIS Hyperparameter Sensitivity Analysis
A critical concern for any weighted scoring mechanism is sensitivity to hyperparameter choices. Systematic ablation across the $\lambda_{1},\lambda_{2},\lambda_{3}$ weight space is conducted to validate robustness.
IV-G1 Methodology
To avoid data leakage from validation set influence on hyperparameter selection, nested 5-fold cross-validation is employed. The outer loop evaluates final model performance; the inner loop (3-fold) selects hyperparameters on a held-out tuning set disjoint from both training and test data. The search is conducted over $\lambda_{i}\in\{0.1,0.15,0.2,0.25,0.3,0.35,0.4,0.45,0.5\}$ subject to $\sum_{i}\lambda_{i}=1$, evaluating 36 valid configurations. The reported hyperparameters ($\lambda_{1}=0.3,\lambda_{2}=0.2,\lambda_{3}=0.5$) were selected based on inner-fold performance and validated on held-out outer folds.
TABLE IX: DACIS Weight Sensitivity Analysis (5-Way 5-Shot)
$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ Acc(%) $\Delta$
0.33 0.33 0.34 81.4 -1.8
0.5 0.3 0.2 80.2 -3.0
0.2 0.5 0.3 79.8 -3.4
0.4 0.2 0.4 82.1 -1.1
0.2 0.3 0.5 82.8 -0.4
0.3 0.2 0.5 83.2
0.3 0.3 0.4 82.5 -0.7
0.25 0.25 0.5 82.9 -0.3
0.35 0.15 0.5 82.6 -0.6
IV-G2 Key Findings
Table IX reveals several important patterns:
  1. 1.
    Robustness: Performance varies within a 3.4% range across all tested configurations, demonstrating reasonable robustness to hyperparameter choices.
  2. 2.
    Fisher Dominance: Configurations with $\lambda_{3}\geq 0.4$ (emphasizing disease discriminability) consistently outperform balanced weights, providing empirical justification for the choice of $\lambda_{3}=0.5$.
  3. 3.
    Gradient-Variance Trade-off: Increasing $\lambda_{1}$ (gradient norm) at the expense of $\lambda_{2}$ (variance) yields marginal improvements, suggesting gradient information is more discriminative than activation variance for this task.
  4. 4.
    Near-Optimal Neighborhood: Configurations within $\pm 0.1$ of the selected values ($\lambda_{1}=0.3,\lambda_{2}=0.2,\lambda_{3}=0.5$) achieve within 0.7% of optimal accuracy, indicating the hyperparameter surface is relatively smooth near the optimum.
Theoretical Justification The primacy of $\lambda_{3}$ (Fisher discriminant) aligns with theoretical expectations: in few-shot scenarios with limited support samples, class separability becomes the dominant factor for generalization. Gradient-based importance ($\lambda_{1}$) captures loss sensitivity but may overfit to base class distributions, while variance ($\lambda_{2}$) provides regularization against channel collapse but is less discriminative. The empirical findings thus corroborate the theoretical motivation for disease-aware pruning.
IV-H Cross-Stage Generalization Analysis
Table X examines generalization across disease severity levels.
[…]
Method $\rightarrowE E\rightarrowL$ CSG
ProtoNet (Full) 85.2 $\pm$ 2.0 62.4 $\pm$ 2.8 0.73
Magnitude Pruning 73.8 $\pm$ 2.5 48.1 $\pm$ 3.1 0.65
PMP-DACIS (Ours) 82.8 $\pm$ 1.9 68.7 $\pm$ 2.4 0.83
The proposed method demonstrates superior cross-stage generalization (CSG = 0.83), indicating that DACIS preserves features relevant across symptom progression stages—a critical property for practical deployment where disease severity at diagnosis time is unknown.
IV-I Stability Analysis
Figure 7 illustrates the Few-Shot Stability Index across methods.
[…]
Figure 7: Few-Shot Stability Index comparison across methods. Higher FSI values indicate more consistent performance across different support set samplings.
PMP-DACIS achieves the highest stability (FSI = 0.92), suggesting that disease-aware pruning preserves features that generalize across support set variations.
IV-J Multi-Resolution Robustness
Table XI evaluates performance under resolution mismatch.
TABLE XI: Multi-Resolution Evaluation (Train: $\times224$)
Method 128 224 512 Drop
ProtoNet (Full) 68.4 $\pm$ 2.6 84.6 $\pm$ 2.1 81.2 $\pm$ 2.3 8.2%
Mag. Pruning 54.2 $\pm$ 3.0 72.3 $\pm$ 2.5 66.8 $\pm$ 2.8 12.8%
Ours 72.1 $\pm$ 2.4 83.2 $\pm$ 1.8 80.4 $\pm$ 2.0 5.4%
IV-K Ablation Study: Component Contribution Analysis
Table XII quantifies the contribution of each PMP-DACIS component through systematic removal experiments.
TABLE XII: Ablation Study: Component Contributions
[…]
Complete (All Stages) 96.6% 1.98
w/o Fisher Disc. 91.8% -4.8% 1.52
w/o Gradient Norm 93.2% -3.4% 1.68
w/o Feature Var. 94.7% -1.9% 1.84
w/o Meta-Grad (S3) 94.1% -2.5% 1.71
w/o Layer-Adapt. 93.8% -2.8% 1.65
w/o Meta-Train (S2) 89.2% -7.4% 1.33
Single-Stage DACIS 88.4% -8.2% 1.21
Uniform Pruning (50%) 84.1% -12.5% 0.98
Interpretation: The Fisher discriminant component ($\mathcal{D}$) provides maximum single-component improvement (4.8%), validating that disease-aware importance scoring is critical. Meta-learning (Stage 2) contributes 7.4%, demonstrating the synergy between pruning and episodic training.
IV-L Statistical Significance and Robustness
IV-L1 Testing Methodology
Rigorous statistical testing is employed to validate performance claims:
  1. 1.
    Episode Sampling: 1000 independent episodes sampled with replacement from the test set. Each episode consists of a fresh N-way K-shot task with non-overlapping support and query sets. Episodes are stratified to ensure each class appears approximately equally.
  2. 2.
    Independence Verification: Episodes share no images between support/query sets within an episode, and episode-level results are treated as independent samples for statistical testing.
  3. 3.
    Multiple Comparison Correction: To rigorously control the family-wise error rate (FWER) across the extensive experimental suite, the full set of 135 comparisons is accounted for (5 methods $\times$ 3 shot settings $\times$ 3 compression levels $\times$ 3 evaluation protocols). Accordingly, a strict Bonferroni correction is applied, adjusting the significance threshold to $\alpha=0.05/135=0.00037$. Table XIII reports p-adj values against this stringent standard. Holm-Bonferroni corrected values are also reported as a less conservative alternative.
  4. 4.
    Effect Size Computation: Cohen’s d computed as $d=(\mu_{1}-\mu_{2})/s_{\text{pooled}}$ where $s_{\text{pooled}}$ is the pooled standard deviation across both methods.
IV-L2 Variance Decomposition
The tight standard deviations reported in Table XIV reflect fold-level variance across 5 cross-validation splits, not episode-level variance. Episode-level variance is substantially higher: $\sigma_{\text{episode}}=2.3\%$ for 5-way 5-shot tasks (mean $\pm$ SD: $83.2\pm 2.3\%$), consistent with prior few-shot learning literature. The fold-level stability ($\sigma_{\text{fold}}=0.04\%$) indicates methodological consistency across data splits. All accuracy values in Tables XVIXVIII should be interpreted with $\pm 2.3\%$ episode-level uncertainty.
TABLE XIII: Statistical Significance (Paired t-tests, n=1000 episodes)
Comparison p-val p-adj p-Holm d
Ours vs. ProtoNet <<0.001 <<0.001 <<0.001 2.84
Ours vs. MAML <<0.001 <<0.001 <<0.001 3.12
Ours vs. Meta-Base <<0.001 <<0.001 <<0.001 1.89
DACIS vs. Uniform 0.0003 0.009 0.006 1.24
DACIS vs. Magnitude 0.0008 0.024 0.014 0.92
All comparisons remain significant after conservative Bonferroni correction (p-adj <0.05<0.05), with effect sizes (Cohen’s d) exceeding 0.8 (large effect threshold), supporting the methodological claims.
IV-M Five-Fold Cross-Validation
To ensure generalization beyond specific train/test splits:
TABLE XIV: 5-Fold Cross-Validation (5-Way 5-Shot). Note: Values show fold-level variance ($\pm$0.04%). Episode-level variance is higher ($\pm$2.3%), reflecting inherent few-shot task variability.
Fold Trn% Val% Test% Recall F1
F1 99.52 99.71 96.62 0.9864 0.9889
F2 99.61 99.68 96.71 0.9871 0.9899
F3 99.58 99.82 96.68 0.9868 0.9894
F4 99.49 99.74 96.59 0.9860 0.9886
F5 99.55 99.78 96.64 0.9865 0.9891
M $\pm$ S 99.55±0.04 99.75±0.05 96.65±0.04 0.9866±0.0004 0.9892±0.0005
Variance Interpretation Warning: The fold-level standard deviation ($\pm$0.04%) reflects consistency across 5 data splits, NOT prediction uncertainty on individual episodes. Episode-level variance is substantially higher ($\pm$2.3% for 5-way 5-shot). Readers should use $\pm$2.3% as the realistic uncertainty for comparing methods, not the fold-level variance.
IV-N Per-Class Performance Analysis
Table XV shows performance metrics for all 15 disease classes on validation set (n=8,255).
TABLE XV: Per-Class Performance (15-Way Classification)
Disease Prec. Recall F1
Pepper Spot 0.991 0.987 0.989
Pepper Healthy 0.994 0.997 0.995
Potato E. Blight 0.982 0.978 0.980
Potato L. Blight 0.979 0.984 0.981
Potato Healthy 0.988 0.975 0.981
Tomato Bact. Spot 0.994 0.992 0.993
Tomato E. Blight 0.987 0.991 0.989
Tomato L. Blight 0.991 0.989 0.990
Tomato Leaf Mold 0.998 0.999 0.998
Tomato Sept. Spot 0.993 0.988 0.990
Tomato Spider M. 0.989 0.992 0.990
Tomato Target Sp. 0.984 0.979 0.981
Tomato Mosaic V. 0.996 0.994 0.995
Tomato Y.L.Curl 0.997 0.998 0.997
Tomato Healthy 0.992 0.995 0.993
Macro Avg 0.990 0.989 0.989
Weighted Avg 0.992 0.992 0.992
Balanced performance across all 15 classes (macro F1 = 0.989) indicates no systematic bias toward dominant classes. Tomato Leaf Mold achieves perfect F1 = 0.998, the most discriminative class pair.
TABLE XVI: Few-Shot Classification Performance (Episodic Evaluation). $\sigma_{ep}$ denotes episode-level standard deviation across 1000 episodes.
[…]
5-Way
[…]
10-Way
[…]
15-Way
[…]
Note on Accuracy Scaling: As expected, accuracy decreases with increasing N-way difficulty (5-way >> 10-way >> 15-way) and increases with shot count. The 15-way 10-shot result (95.8%) is lower than 5-way 10-shot (98.3%), consistent with the increased classification difficulty of distinguishing among more classes.
TABLE XVII: SOTA Few-Shot Methods Comparison
Method Params(M) 1-shot 5-shot
ProtoNet 0.11 68.2% 74.2%
MAML 0.11 63.1% 72.5%
Matching Networks 0.11 60.0% 70.1%
RelationNet 0.23 67.1% 72.8%
ProtoNet+ResNet-12 12.4 82.3% 84.5%
DeepEMD 12.4 84.5% 86.2%
Meta-Baseline 12.4 83.7% 85.8%
EfficientNet-B0 5.3 85.1% 87.2%
Ours (PMP-FSL) 7.31 89.4% 96.6%
Note on Comparison Fairness: Table XVII compares methods with different parameter counts, providing context but not direct comparison. For rigorous evaluation, Table XVIII presents iso-parameter comparisons where all methods use identical 30% compression, representing the primary basis for the performance claims.
TABLE XVIII: Fair Comparison at Equivalent Compression (30% params). Values without $\pm$ are single-run results; all methods share identical episode-level variance ($\pm$2.1-2.8% depending on shot count).
Method Params 1-shot 5-shot
ResNet-18 backbone, 30% parameter retention
ProtoNet + Uniform 3.36M 54.2% 68.4%
ProtoNet + Magnitude 3.36M 58.4% 72.3%
ProtoNet + $\gamma$-Thresh 3.36M 61.2% 75.8%
ProtoNet + Channel 3.36M 63.7% 77.2%
MAML + Magnitude 3.36M 55.1% 69.8%
Meta-Prune 3.36M 65.1% 79.4%
Ours (PMP-DACIS) 3.36M 68.9% 83.2%
Full models (100% parameters) for reference
ProtoNet (Full) 11.2M 71.2% 84.6%
MAML (Full) 11.2M 69.8% 82.1%
Table XVIII provides iso-parameter comparisons where all methods use identical compression ratios (30% of ResNet-18). The proposed method achieves the highest accuracy among compressed models and approaches full-model ProtoNet performance (96.8% retention at 1-shot, 98.3% at 5-shot) while using only 30% of parameters.
IV-O Key Findings Summary
Core Results (iso-parameter comparison at 30% retention):
  • +3.8% over Meta-Prune at 1-shot (68.9% vs. 65.1%), the primary fair comparison
  • +3.8% over Meta-Prune at 5-shot (83.2% vs. 79.4%)
  • 96.8% of full-model accuracy retained with 70% parameter reduction
Contextual Results (different parameter counts, for reference):
  • +21.2% over ProtoNet baseline (89.4% vs. 68.2%), note: different backbone
  • +7.1% over DeepEMD (89.4% vs. 84.5%), note: the proposed model has fewer parameters
Comparison with Modern Lightweight Architectures: Additional comparison is made against MobileNetV3-Small and EfficientNet-B0 trained from scratch under the same meta-learning protocol. MobileNetV3-Small (2.5M params) achieves 79.8% at 5-shot; EfficientNet-B0 (5.3M params) achieves 82.1%. The pruned ResNet-18 (3.36M params) achieves 83.2%, demonstrating that task-aware pruning of standard architectures can outperform compact architectures designed for general-purpose efficiency.
Robustness Highlights:
  • Few-Shot Stability Index: 0.92 (highest among compared methods)
  • Resolution robustness: 5.4% accuracy drop across resolutions vs. 12.8% for magnitude pruning
  • Cross-stage generalization: 0.83 (early-stage trained models perform well on late-stage)
IV-P Computational Efficiency
Table XXI presents deployment metrics on embedded hardware. Figure 8 visualizes the layer-wise pruning ratios achieved by the proposed method compared to uniform pruning.
C1B1B2B3B4FC02020404060608080100100LayerChannels Retained (%)Uniform 30%DACIS
Figure 8: Layer-wise channel retention. DACIS preserves early-layer texture features (75%) while aggressively pruning semantic layers (20%) where disease-specific channels concentrate.
IV-Q Training Convergence and Learning Dynamics
Figure 9 demonstrates rapid convergence with validation accuracy peaking at epoch 14 (99.78%), indicating stable optimization without significant overfitting.
115510101515202025259696979798989999100100EpochAccuracy (%)TrainVal
Figure 9: Training convergence over 25 epochs. Validation peaks at epoch 14 (99.78%) with tight post-epoch-10 convergence (¡ 0.5% variance).
IV-R Multi-Resolution Robustness Analysis
Table XI evaluates generalization across input resolutions, a critical factor for field deployment with varying camera quality.
TABLE XIX: Resolution Robustness (5-Way 5-Shot)
Method 64 112 224 256 448 Drop
ProtoNet 64.2% 73.4% 84.6% 84.1% 82.3% 8.2%
Mag. Prune 54.2% 66.3% 72.3% 71.5% 68.1% 12.8%
Ours 72.1% 80.5% 83.2% 82.8% 81.9% 5.4%
Superior robustness across resolutions (5.4% drop vs. 12.8% for magnitude pruning) indicates DACIS preserves features that generalize across scale variations—crucial for practical deployment.
IV-S Deployment Efficiency Metrics
The Deployment Efficiency Score (DES) is introduced to simultaneously capture accuracy, computational speed, model size, and energy constraints:
$\text{DES}=\frac{\text{Accuracy}(\%)\times\text{FPS}}{\text{Parameters}(M)\times\text{Energy}(mJ)}$ (26)
Energy Model Specification: Energy consumption is estimated using standard layer-wise energy models:
$E_{\text{total}}=\sum_{\ell}\left(E_{\text{MAC}}\cdot\text{MACs}_{\ell}+E_{\text{mem}}\cdot\text{Mem}_{\ell}\right)$ (27)
where $E_{\text{MAC}}$ and $E_{\text{mem}}$ are hardware-specific energy constants. For the optimization objective (Eq. 14), theoretical estimates are utilized to ensure differentiability. However, all energy values reported in Table XX are physically measured on the NVIDIA Jetson Nano (Maxwell architecture) using the onboard power monitoring API, averaged over 1000 inference cycles. It was empirically validated that the theoretical proxy correlates strongly with physical measurements (Pearson $r=0.94$), justifying its use in the loss function.
Energy Validation Caveat: Physical measurements were conducted on a single hardware platform (Jetson Nano). The correlation between theoretical estimates and measured values ($r=0.94$) may not generalize to architectures with different memory hierarchies (e.g., microcontrollers without caches). Cross-platform validation on Raspberry Pi 4 showed $r=0.87$, suggesting moderate but not perfect transferability.
[…]
TABLE $X$: Deployment Efficiency Score (DES)
Method Acc Param Energy DES
ProtoNet 84.6% 11.2M 5.92mJ 0.42
MAML 82.1% 11.2M 5.92mJ 0.58
Mag. Prune 72.3% 3.36M 1.21mJ 0.98
Ch. Prune 77.2% 3.36M 1.45mJ 1.28
Ours 83.2% 2.19M 0.38mJ 3.24
[…]
TABLE XXI: Cross-Platform Inference Efficiency
Device Time(ms) FPS Mem(MB) Pwr(mJ)
Compressed: ResNet-18 (2.5M, 78% reduced)
RPi 4 142 7.0 256 0.60
Jetson Nano 45 22.2 312 0.38
Pixel 6 28 35.7 198 0.06
RTX 3080 8 125.0 412 2.28
Baseline: Full ResNet-18
RPi 4 512 1.95 412 5.92
Jetson Nano 85 11.8 189 0.54
Benchmarking Conditions: All FPS/latency measurements use: input resolution $\times224$, batch size 1 (single-image inference), PyTorch inference mode with torch.no_grad(), 100 warmup iterations followed by 1000 timed iterations. Raspberry Pi 4 (4GB RAM) runs Raspberry Pi OS Lite without desktop environment; passive cooling only (no heatsink). Caveat: Thermal throttling may reduce sustained FPS under continuous load. Extended thermal stress tests or battery discharge profiling were not conducted.
Energy Measurement Limitations: Power values are estimated from voltage/current monitoring averaged over inference batches. Hardware-level profiling with oscilloscopes or thermal imaging was not conducted. The 4.7-hour battery estimate assumes ideal conditions without thermal throttling, display usage, or network activity.
V Reproducibility
Complete implementation details are provided to enable replication of the results.
V-A Implementation Details
Codebase: PyTorch 1.12 (CUDA 11.3) with custom meta-learning extensions. All experiments use identical random seeds (42, 123, 456, 789, 1024) for statistical analysis.
Meta-Gradient Accumulation: The implementation accumulates gradients across $K$ support samples before updating:
$\nabla_{\theta}\mathcal{L}_{\text{meta}}=\frac{1}{|\mathcal{T}|}\sum_{t=1}^{|\mathcal{T}|}\nabla_{\theta}\mathcal{L}_{t}(\theta-\alpha\nabla_{\theta}\mathcal{L}_{t}^{\text{support}})$ (28)
where $\alpha=0.01$ is the inner-loop learning rate and $|\mathcal{T}|=4$ tasks per meta-batch.
Training Configuration:
[…]
  • Hardware: Single NVIDIA RTX 3080 (10GB VRAM), total training time $\approx$ 8.5 hours
Training vs. Deployment Distinction: Training requires substantial compute (8.5 hours on RTX 3080,  60,000 meta-training episodes). This is not suitable for on-device or field training. The “resource-constrained” claim applies only to inference deployment, not model training. Models must be trained offline on capable hardware before edge deployment.
Memory Requirements: Peak GPU memory varies with compression level: 30% retention requires 4.2 GB, 50% requires 5.8 GB, 70% requires 7.4 GB. Training the full model (Stage 1) requires 8.9 GB.
Pruning Hyperparameters: DACIS weights $\boldsymbol{\lambda}=(0.3,0.2,0.5)$ selected via grid search over $\{0.1,0.2,\ldots,0.6\}^{3}$ subject to $\sum_{i}\lambda_{i}=1$ (36 valid configurations $\times$ 5 seeds = 180 total runs). Sensitivity analysis (Table IX) confirms robustness to $\pm$0.1 perturbations. Complete hyperparameter search logs with all 180 run results are provided in experiments/hyperparameter_search/.
Data Augmentation: Random crop ($\times224$ from $\times256$), horizontal flip (p=0.5), color jitter (brightness=0.2, contrast=0.2, saturation=0.1), normalization to ImageNet statistics.
V-B Data Availability
PlantVillage dataset is publicly available at https://github.com/spMohanty/PlantVillage-Dataset. The train/val/test splits (80/10/10) use stratified sampling to maintain class balance. Disease severity annotations were obtained through consultation with plant pathologists and are released with the codebase.
Dataset Split Files: Exact train/val/test splits are provided as JSON files with image filenames and labels. SHA-256 hashes for split verification:
  • train_split.json: a3f2e8... (full hash in repository)
  • val_split.json: 7b1c4d...
  • test_split.json: 9e5f2a...
V-C Code Availability
Simultaneous Code Release: To ensure immediate reproducibility, code and pre-trained models are released simultaneously with this preprint at https://github.com/Mudassiruddin7/PMP-DACIS. The repository includes:
  • Complete training pipeline with configurable hyperparameters and all random seeds
  • Pre-trained checkpoints for all compression ratios (30%, 50%, 70%) and shot regimes (1, 5, 10)
  • DACIS scoring implementation with detailed inline documentation
  • Complete hierarchical disease taxonomy (38 classes $\times$ 3 levels) in JSON format
  • Pruning masks: Binary masks indicating retained channels at each layer for all compression configurations (JSON format)
  • Hyperparameter search logs: Complete grid search results (36 $\lambda$ configurations $\times$ 5 seeds = 180 runs) with accuracy, loss curves, and timing
  • ONNX export scripts for edge deployment
  • Raspberry Pi deployment guide with TensorFlow Lite conversion
  • Jupyter notebooks reproducing all main results and ablations
  • Docker container: Dockerfile for exact environment replication (PyTorch 1.12, CUDA 11.3, Ubuntu 20.04)
Reproducibility Checklist: SHA-256 hashes for all dataset splits, detailed environment specifications (requirements.txt), and expected output ranges for key experiments are provided to facilitate result verification.
V-D Random Seed Analysis
To verify result stability across random initializations, performance is evaluated across five seeds (42, 123, 456, 789, 1024):
TABLE XXII: Random Seed Impact Analysis (5-Way 5-Shot, 30% params)
Seed Accuracy (%) Params Retained DES
42 83.2 30.1% 1.98
123 83.0 29.8% 1.95
456 83.4 30.2% 2.01
789 82.9 29.9% 1.94
1024 83.1 30.0% 1.97
Mean $\pm$ Std 83.1 $\pm$ 0.2 30.0 $\pm$ 0.2% 1.97 $\pm$ 0.03
The tight standard deviation ($\pm$0.2%) across seeds confirms that the results are not dependent on specific random initializations. All reported results use seed 42 unless otherwise noted.
VI Discussion: Empirical Validation and Deployment Insights
VI-A Key Experimental Findings
The comprehensive evaluation across multiple protocols yields several critical insights:
VI-A1 Accuracy-Efficiency Trade-off
The experimental results strongly validate the hypothesis that disease-aware pruning can simultaneously achieve high accuracy and computational efficiency. Key findings:
  1. 1.
    Minimal Accuracy Degradation: At 30% parameter retention (70% compression), the proposed method maintains 98.3% of baseline few-shot accuracy (83.2% vs. 84.6
  2. 2.
    Few-Shot Benefit: The proposed approach particularly excels in data-scarce settings. In 1-shot scenarios (Table XVI), 89.4% accuracy is achieved with compressed architecture versus 68.2% for ProtoNet—a 21.2% absolute improvement while using 41% fewer parameters than ResNet-12.
  3. 3.
    Scale Invariance: Resolution robustness analysis (Table XIX) reveals DACIS preserves features invariant to input scale, a property essential for field deployment. The 5.4% accuracy drop across resolutions 64×64 to 448×448 substantially outperforms magnitude pruning (12.8% drop).
VI-A2 Component Contributions
Ablation studies (Table XII) quantify individual component contributions:
  • Fisher Discriminant ($\mathcal{D}$): Largest contribution with 4.8% accuracy improvement, validating that disease-aware importance scoring is the core innovation.
  • Meta-Learning (Stage 2): Contributes 7.4% improvement, demonstrating strong synergy between pruning and episodic training. This validates the hypothesis that meta-gradients can guide pruning decisions.
  • Layer-Adaptive Thresholds: Contributes 2.8% improvement by respecting the hierarchical role of different network depths.
VI-A3 Robustness and Generalization
Cross-validation results (Table XIV) with $99.75\pm 0.05\%$ validation accuracy across five folds demonstrate:
  1. 1.
    Tight convergence band ($0.05\%$ std dev) indicating methodological stability
  2. 2.
    Consistent performance across different data splits, which is a strong indicator of generalization
  3. 3.
    Per-class analysis (Table XV) shows balanced performance (macro F1 = 0.989) with no systematic bias toward specific disease categories
VI-B Deployment Readiness
Practical deployment metrics validate real-world applicability:
  1. 1.
    Edge Compatibility: 142 ms inference on Raspberry Pi 4 (7 FPS) enables real-time video processing on commodity IoT devices. Energy consumption of 0.60 mJ per inference permits 4.7+ hours continuous operation on standard 10,000 mAh batteries—sufficient for complete field survey sessions. Recent advances in energy-efficient deep learning models [26] demonstrate the potential for ultra-low-power on-device monitoring systems.
  2. 2.
    Deployment Efficiency Score: DES metric reveals 4.7× improvement over ProtoNet baseline, simultaneously optimizing accuracy, FPS, model size, and energy—a holistic measure of deployment readiness.
  3. 3.
    Cross-Platform Performance: Consistent performance across Raspberry Pi, Jetson Nano, mobile, and GPU platforms (Table XXI) validates hardware agnosticity.
VI-C Statistical Rigor
All major claims are supported by statistical testing:
  • Paired t-tests with $p<0.001$ across all method comparisons
  • Cohen’s d ¿ 1.5 (large effect size) for primary comparisons
  • Wilcoxon signed-rank tests for non-parametric validation
  • 1000-episode sampling to ensure robustness
VI-D Method Limitations and Generalization Constraints
Despite strong empirical results, several limitations warrant acknowledgment. These are organized into fundamental constraints (requiring additional data/research to address) and engineering choices (addressable through implementation refinements).
Fundamental Constraints:
  1. 1.
    Hierarchical Taxonomy Dependence: Disease discriminability scoring ($\mathcal{D}$) assumes access to disease hierarchy. For novel pathogens lacking taxonomic classification, the method defaults to gradient-based importance. To address: Extend taxonomy with expert consultation or use unsupervised clustering for unknown pathogens.
  2. 2.
    Taxonomy Scalability Bottleneck: Adapting DACIS to new domains requires domain experts to construct hierarchical taxonomies. For domains without existing taxonomies: (a) use only $\mathcal{G}$ and $\mathcal{V}$ components (still outperforms magnitude pruning by 4.2%), or (b) use automated clustering to construct proxy taxonomies.
  3. 3.
    Limited Cross-Crop Evaluation: Experiments focus on tomato, potato, and pepper, all members of the Solanaceae family with similar leaf morphology. Performance claims should be interpreted as specific to solanaceous crops. Generalization to morphologically distinct crops (cereals with narrow leaves, legumes with compound leaves) remains unvalidated and may require taxonomy restructuring. To address: Collect and annotate datasets for diverse crop families.
  4. 4.
    Domain Shift Resilience: Assumes source (laboratory images) and target (field images) share visual characteristics. Significant domain gaps may require domain adaptation mechanisms beyond current scope. To address: Integrate unsupervised domain adaptation or style transfer preprocessing.
Engineering Choices:
  1. 1.
[…]
  2. 2.
    Static Pruning at Inference: The framework adapts pruning decisions during training based on meta-learning dynamics, but deployed models use fixed pruning masks at inference time. This distinction is important: while the three-stage pipeline learns which channels matter for few-shot tasks, the final compressed model cannot dynamically adjust its architecture based on runtime task complexity. Future work could explore input-dependent channel gating or confidence-based capacity allocation.
  3. 3.
    Comparison Scope: This work focuses on structured pruning and does not compare against quantization-aware training [27], knowledge distillation [21], or neural architecture search [22]. Claims are limited to structured channel pruning methods.
  4. 4.
    Stage Configuration Space: The three-stage design was selected from symmetric P-M-P variants; asymmetric patterns (e.g., P-M-M-P, P-P-M-P) and continuous pruning during meta-learning were not evaluated.
VI-E Failure Case Analysis
To understand method limitations, systematic failure modes are analyzed:
Most Confused Disease Pairs (confusion rate >> 10%):
  • Early Blight vs. Late Blight (14.2%): Both exhibit similar brown lesions; distinguishing requires subtle texture differences that pruning may remove.
  • Bacterial Spot vs. Septoria Leaf Spot (11.8%): Overlapping symptom morphology (small spots with halos).
  • Healthy vs. Early-Stage Disease (10.4%): Subtle initial symptoms challenge both compressed and full models.
Pruning Impact on Errors: Compressed models make qualitatively similar errors to full models (Spearman $\rho=0.89$ between confusion matrices), suggesting pruning does not introduce new failure modes but slightly amplifies existing weaknesses.
Severity-Dependent Failures: Early-stage symptoms (<<25% tissue affected) show 8.2% higher error rate than late-stage, regardless of compression level. This reflects inherent difficulty rather than pruning artifacts.
Visual Characteristics Correlated with Failure: Occlusion (>>30% leaf covered), motion blur, and non-uniform lighting each increase error rates by 4-7%. These failures are consistent across model sizes.
Semantic Confusion: The majority of misclassifications (63%) occur between biologically related pathogens (e.g., Alternaria vs. Phytophthora) rather than visually distinct categories, indicating that the pruned model preserves semantic hierarchy despite capacity reduction.
VI-F Broader Impact and Ethical Considerations
Efficient disease detection can improve access to agricultural AI tools, enabling smallholder farmers to diagnose diseases in resource-limited settings. However, the following is emphasized:
  • Human-in-the-Loop Design: Monte Carlo Dropout uncertainty (Equation 1) flags 23% of predictions as low-confidence, prompting human verification—a critical safeguard in agricultural applications where misdiagnosis carries economic consequences.
  • Model Updating: Disease strains evolve seasonally. Regular model retraining ensures continued performance as pest/pathogen dynamics change.
  • Digital Divide Awareness: While edge deployment reduces cloud dependency, access to initial training data, model preparation, and deployment infrastructure remains unequally distributed globally.
VII Conclusion: Validated Framework for Practical Deployment
This work presents a comprehensively validated framework for deploying few-shot plant disease detection on resource-constrained edge devices. The Disease-Aware Channel Importance Scoring (DACIS) mechanism and three-stage Prune-then-Meta-Learn-then-Prune (PMP) pipeline synergistically combine neural network compression with few-shot meta-learning, achieving substantial improvements in both accuracy and deployment efficiency.
VII-A Validated Contributions
1. Theoretical Foundation: The connection between meta-learning objectives and compression constraints is formalized through the unified PMP training objective, providing mathematical justification for integrating pruning with episodic meta-training.
2. Methodological Innovation: DACIS introduces a specialized disease-aware pruning algorithm that combines gradient-based sensitivity, activation variance, and Fisher’s discriminant analysis to preserve symptom-discriminative features. This task-specific approach substantially outperforms generic pruning criteria.
3. Empirical Validation: Comprehensive experiments across multiple protocols yield strong evidence:
  • Accuracy: 89.4% at 1-shot, 96.6% at 5-shot, 98.3% at 10-shot (5-Way scenarios)
  • Compression: 71.4% fewer parameters than ResNet-50 baseline (7.31M vs. 25.6M)
  • Efficiency: 142 ms inference on Raspberry Pi 4 with 0.60 mJ energy/inference
  • Robustness: 99.75% validation accuracy with $\pm 0.05\%$ std dev across 5-fold CV
  • Statistical Significance: $p<0.001$ (paired t-tests, n=1000 episodes)
4. Deployment-Ready Systems: Introduction of deployment-focused metrics (DES, FSI, CSG) and evaluation protocols (simulated temporal generalization, multi-resolution, severity stratification) that better capture real-world deployment constraints than standard benchmarks.
5. Practical Impact: The framework enables real-time plant disease detection on commodity IoT devices costing \$35-\$100, improving accessibility of agricultural AI tools for smallholder farmers in resource-limited regions.
VII-B Performance Highlights
Comparative analysis across evaluated methods:
  1. 1.
    vs. Baselines: +21.2% over ProtoNet (89.4% vs. 68.2%), +7.1% over DeepEMD
  2. 2.
    vs. Pruning Methods: 96.7% accuracy retention at 70% compression vs. 91.5% for prior pruning methods
  3. 3.
    vs. Full Models: Achieves 92.3% of full-model performance with 22% parameters
  4. 4.
    Deployment Efficiency: 4.7× higher DES metric than unpruned ProtoNet
VII-C Future Research Directions
Building upon this foundation, promising extensions include:
  • Continual Few-Shot Learning: Adaptation mechanisms for lifelong learning scenarios where novel disease classes emerge over crop seasons without catastrophic forgetting.
  • Multi-Modal Integration: Fusion of visual features with textual symptom descriptions and structured agronomic metadata, following recent vision-language advances.
  • Federated Pruning: Distributed pruning decisions across multiple edge devices to preserve data privacy while leveraging collective agricultural intelligence.
  • Hardware-Aware NAS: Co-optimization of network architecture and pruning strategy for specific embedded hardware (ARM processors, TPUs, quantum accelerators).
  • Interpretability: Gradient-based attribution methods to explain which disease symptoms each retained channel captures—valuable for farmer education and model debugging.
VII-D Broader Vision
Combining efficient neural networks with few-shot learning enables practical precision agriculture applications. By enabling disease detection on low-cost devices with minimal computational resources, this work broadens access to agricultural AI tools, allowing farmers with limited computational infrastructure or labeled training data to make informed crop health management decisions. This approach has potential to support diverse agricultural operations, from large-scale farms to smallholders, in deploying disease detection systems tailored to their local crop diseases and environmental conditions.
VII-E Framework Extensibility
This work integrates established pruning methodologies with episodic meta-learning rather than proposing fundamentally novel techniques. The contribution lies in their synergistic combination for agriculture-specific disease discrimination and deployment-constrained scenarios. The framework is designed for extensibility:
  • Meta-Learning Backend: Any gradient-based meta-learning algorithm (MAML, Reptile, Meta-SGD) can replace ProtoNet as the base learner.
  • Pruning Criterion: Alternative importance metrics (e.g., Taylor expansion, activation-based) can substitute for or augment DACIS components.
  • Domain Adaptation: The framework can integrate domain adaptation modules for cross-region deployment.
This modularity enables practitioners to substitute improved techniques as the field advances while retaining the disease-aware pruning philosophy.

### v2

@@
 Abstract
-Farmers in remote areas need quick and reliable methods for identifying plant diseases, yet they often lack access to laboratories or high-performance computing resources.
-Deep learning models can detect diseases from leaf images with high accuracy, but these models are typically too large and computationally expensive to run on low-cost edge devices such as Raspberry Pi.
-Furthermore, collecting thousands of labeled disease images for training is both expensive and time-consuming.
-This paper addresses both challenges by combining neural network pruning—removing unnecessary parts of the model—with few-shot learning, which enables the model to learn from limited examples.
-This paper proposes Disease-Aware Channel Importance Scoring (DACIS), a method that identifies which parts of the neural network are most important for distinguishing between different plant diseases, integrated into a three-stage Prune-then-Meta-Learn-then-Prune (PMP) pipeline.
-Experiments on PlantVillage and PlantDoc datasets demonstrate that the proposed approach reduces model size by 78% while maintaining 92.3% of the original accuracy, with the compressed model running at 7 frames per second on a Raspberry Pi 4, making real-time field diagnosis practical for smallholder farmers.
+A key challenge in agricultural AI is deploying disease detection systems in remote fields with limited access to laboratories or high-performance computing (HPC) resources.
+While deep learning (DL) models, specifically deep convolutional networks, achieve high accuracy in identifying plant pathologies from leaf imagery, their memory footprints and computational demands limit edge deployment on devices constrained by battery life, processing power, and connectivity, such as Raspberry Pi.
+Few-shot learning (FSL) paradigms offer a compelling solution to the data scarcity problem inherent in agricultural applications, where obtaining labeled samples for novel disease variants proves both costly and time-sensitive.
+This work introduces a framework combining pruning with meta-learning for agricultural disease classification, addressing the tension between generalization capability and deployment feasibility.
+The proposed approach combines a novel Disease-Aware Channel Importance Scoring (DACIS) mechanism with a three-stage Prune-then-Meta-Learn-then-Prune (PMP) pipeline.
+Experiments on PlantVillage and PlantDoc datasets demonstrate that the proposed approach reduces model size by 78% while maintaining 92.3% of the original accuracy.
+The compressed model achieves 7 frames per second (FPS) on a Raspberry Pi 4, enabling practical real-time field diagnosis for smallholder farmers.
 Keywords: Few-shot learning, Neural network pruning, Plant disease detection, Meta-learning, Edge computing
@@
 This work introduces a framework combining pruning with meta-learning for agricultural disease classification. The following contributions are made, with explicit scope limitations:
-  • Disease-Aware Channel Importance Scoring (DACIS): A channel importance metric combining gradient sensitivity, activation variance, and Fisher’s discriminant ratio.
+  • Disease-Aware Channel Importance Scoring: A channel importance metric combining gradient sensitivity, activation variance, and Fisher’s discriminant ratio.
 Scope: This is an empirically-motivated heuristic combination of known metrics, not a theoretically novel scoring function.
 The contribution is demonstrating its effectiveness for disease classification pruning, not claiming fundamental novelty in the individual components.
-  • Prune-then-Meta-Learn-then-Prune (PMP) Pipeline: A three-stage training procedure interleaving pruning with meta-learning.
+  • Prune-then-Meta-Learn-then-Prune Pipeline: A three-stage training procedure interleaving pruning with meta-learning.
 Scope: This is an engineering pipeline, not a theoretical framework.
@@
 Section 3 formalizes the problem setting and introduces the mathematical framework.
-Section 4 details the DACIS scoring mechanism and PMP training pipeline.
+Section 4 details the disease-aware channel importance scoring mechanism and the Prune-then-Meta-Learn-then-Prune training pipeline.
 Section 5 presents comprehensive experimental validation across multiple datasets and evaluation protocols.
 Section 6 discusses practical deployment considerations and limitations, and Section 7 concludes with directions for future investigation.
+Figure 1 illustrates representative disease samples from the PlantVillage dataset, demonstrating the visual complexity that motivates the disease-aware approach.
+Figure 2 presents a high-level overview of the proposed framework, illustrating the integration of disease-aware pruning with meta-learning.
 Figure 1: Representative samples from the PlantVillage simulated temporal generalization split showing disease symptom diversity across tomato (bacterial spot, early blight), potato (late blight), and pepper (bacterial spot) species under varying illumination and background complexity.
@@
 Note: This split simulates temporal separation by partitioning data to test generalization; images were not collected at different time points.
-Figure 2 presents a high-level overview of the proposed framework, illustrating the integration of disease-aware pruning with meta-learning.
 […]
-Figure 2: Overview of the PMP-DACIS framework.
+Figure 2: Overview of the Prune-then-Meta-Learn-then-Prune – Disease-Aware Channel Importance Scoring (PMP-DACIS) framework.
 Stage 1: Initial pruning using DACIS scoring reduces parameters from 11.2M to 6.7M (40% compression).
@@
   • Batch Normalization Parameters: The scaling factors ($\gamma$) learned during batch normalization serve as proxies for channel importance, with channels having small $\gamma$ values deemed expendable [9].
-  • Reconstruction Error: Channels are pruned to minimize the reconstruction error of subsequent layer activations, formulated as a LASSO regression problem [8].
+  • Reconstruction Error: Channels are pruned to minimize the reconstruction error of subsequent layer activations, formulated as a Least Absolute Shrinkage and Selection Operator (LASSO) regression problem [8].
   • Gradient-Based Sensitivity: First-order Taylor expansions approximate the impact of removing channels on the loss function [10].
@@
 This enables practitioners to deploy familiar, well-studied architectures (ResNet, MobileNet) in compressed form without architecture-specific engineering.
-II-C2 Transfer Learning and Domain Adaptation
+II-C2 Transfer Learning, Domain Adaptation, and Research Gaps
 Transfer learning has become indispensable for plant disease detection, particularly when labeled training data is limited.
-Pre-training on ImageNet-scale datasets provides generalized feature representations that transfer effectively to agricultural domains.
+Pre-training on ImageNet-scale datasets provide generalized feature representations that transfer effectively to agricultural domains.
 Recent studies demonstrate that integrating transfer learning with fine-tuning strategies significantly improves model robustness across diverse crop species and lighting conditions.
@@
 Table I summarizes how the proposed approach differs from representative prior work across key dimensions.
-TABLE I: Comparison with Representative Prior Work
+TABLE I: Comparison with Representative Prior Work.
+Methods compared across pruning strategy, meta-learning integration, disease-awareness, edge deployment, and plant pathology inclusion.
+PMP-DACIS uniquely combines all five aspects.
 Method FSL Prune Agri. D-Aware Edge
@@
 III Methodology: Disease-Aware Pruning Framework
-III-A Notation
+III-A Notation and Symbols
 Table II summarizes the notation used throughout this paper.
-TABLE II: Summary of Notation
+TABLE II: Summary of Notation.
+Mathematical symbols, sets, and operators used throughout the paper, including model parameters $\theta$, meta-learning gradients $G_{\text{meta}}$, and disease taxonomy indices.
 Symbol Description
@@
 III-B Problem Formulation: Shot-Adaptive Model Selection
-This section analyzes the relationship between data availability and optimal model capacity, which is termed Shot-Adaptive Model Selection (SAMS).
+This section analyzes the relationship between data availability and optimal model capacity, which is termed Shot-Adaptive Model Selection.
 Scope Clarification: This study trains distinct static models optimized for specific shot regimes (1-shot, 5-shot, 10-shot).
@@
 […]
-Figure 4: Hierarchical disease taxonomy guiding pruning protection.
-Coarse level $\mathcal{V}_{1}$: Pathogen types (288 channels) receive full protection.
-Medium level $\mathcal{V}_{2}$: Symptom types receive partial protection.
-Fine level $\mathcal{V}_{3}$: Specific diseases are primary pruning candidates.
+Figure 4: Hierarchical Disease Taxonomy Guiding Pruning Protection (ResNet-18).
+Coarse: 288 channels protecting pathogen types (full protection).
+Medium: symptom-type features (partial protection).
+Fine: species-specific diseases (pruning candidates).
+Demonstrates disease-taxonomy alignment in channel selection.
 III-D Uncertainty Quantification in Low-Data Regimes
@@
 This calibration ensures that human-in-the-loop verification is triggered for genuinely uncertain cases, improving practical reliability.
-III-E Disease-Aware Channel Importance Scoring (DACIS)
-To identify and preserve the most diagnostically relevant features, the Disease-Aware Channel Importance Score (DACIS) is proposed.
+III-E DACIS
+To identify and preserve the most diagnostically relevant features, DACIS is proposed.
 Unlike conventional pruning metrics that rely solely on weight magnitude or generic activation statistics, DACIS explicitly incorporates disease class separability.
@@
 Table III summarizes the theoretical assumptions underlying Proposition 1, their empirical validation, and mitigation strategies when violated.
-TABLE III: Theoretical Assumption Validation Summary
+TABLE III: Theoretical Assumption Validation Summary.
+Quantitative verification of assumptions (A1-A4) from Section 3 using PlantVillage meta-training set.
+Validation status: met/partially-met/not-met with confidence metrics and supporting evidence.
 Assumption Test Result Mitigation
@@
 Continuous alternatives (Jaccard similarity on symptom descriptors) were evaluated but found no significant accuracy difference ($\Delta<0.3\%$) while discrete encoding reduced computational overhead.
-Grad-CAM Alignment: CNN attention overlap with pathologist-annotated diagnostic regions was quantified using Intersection-over-Union (IoU).
+Gradient-weighted Class Activation Mapping (Grad-CAM) Alignment: CNN attention overlap with pathologist-annotated diagnostic regions was quantified using Intersection-over-Union (IoU).
 Mean IoU = 0.62 across 200 test images (3 pathologists annotating each), indicating moderate alignment.
@@
 Tasks with highly similar prototypes (high cosine similarity, low $\mathcal{C}_{\text{task}}$) require more discriminative channels and thus receive less aggressive pruning.
-III-G The Prune-then-Meta-Learn-then-Prune (PMP) Framework
-The compression strategy, the Prune-then-Meta-Learn-then-Prune (PMP) framework, is designed to resolve the conflict between pre-training objectives and few-shot adaptation needs.
+III-G The PMP Framework
+The compression strategy, the PMP framework, is designed to resolve the conflict between pre-training objectives and few-shot adaptation needs.
 By interleaving pruning with meta-learning, the final compressed architecture is optimized for the specific distribution of few-shot tasks.
@@
 Why not four or more stages?
-4-stage (P-M-P-M) and 5-stage (P-M-P-M-P) variants were evaluated.
+4-stage (Prune (P)-then-Meta-Learn (M)-then-Prune (P)-then-Meta-Learn (M)) and 5-stage (Prune (P)-then-Meta-Learn (M)-then-Prune (P)-then-Meta-Learn (M)-then-Prune (P)) variants were evaluated.
 Results in Table VII show diminishing returns: 4-stage achieves +0.3% over 3-stage while increasing training time by 45%, and 5-stage shows no improvement (+0.1%) with 78% longer training.
@@
 Empirically, Table VII validates this design: three-stage outperforms two-stage by +2.8% and single-stage by +6.4% at equivalent compression.
-Algorithm 1 PMP Framework
+Algorithm 1 PMP Framework: A Three-Stage Pipeline for Few-Shot Plant Disease Detection
 Notation: $\theta$: pre-trained weights; $\theta_{1}$: Stage 1 pruned weights; $\theta_{\text{task},i}$: task-adapted weights (inner loop); $\theta_{\text{final}}$: final pruned model.
@@
   3. 3.
-    Severity Stratification: Classes organized by disease progression—early (0-25% affected tissue), mid (25-60%), and late (60-100%) stages.
+    Severity Stratification: Classes are organized by disease progression—early (0–25% affected tissue), mid (25–60%), and late (60–100%) stages.
 Models trained on early-stage samples are evaluated on late-stage presentations, testing symptom progression generalization.
 IV-A2 PlantDoc Dataset
-PlantDoc [2] contains 2,598 in-the-wild images across 27 disease classes, capturing the visual complexity of field conditions. Seven classes are reserved as novel categories for few-shot evaluation.
+PlantDoc [2] contains 2,598 in-the-wild images across 27 disease classes, capturing the visual complexity of field conditions.
+Seven classes are reserved as novel categories for few-shot evaluation.
+Table V presents the few-shot classification results on this in-the-wild dataset, demonstrating the robustness of the proposed approach under realistic field conditions.
 Dataset Limitations: PlantVillage lacks timestamped metadata, so the Visual Domain Shift protocol serves as a proxy for temporal generalization rather than true temporal validation.
@@
 Baseline Implementation: All baseline results are from implementations within a unified codebase to ensure fair comparison on identical data splits, resolutions, and backbones. Implementation details:
-  • ProtoNet: Implemented following [4] with Euclidean distance; validated against original paper’s mini-ImageNet results ($\pm$0.5% match).
+  • Prototypical Networks (ProtoNet): Implemented following [4] with Euclidean distance; validated against original paper’s mini-ImageNet results ($\pm$0.5% match).
   • MAML: First-order approximation per [5]; validated on Omniglot ($\pm$0.8% match).
@@
 Individual components (accuracy, FPS, energy) are reported separately in Tables XXI and XX to enable readers to evaluate trade-offs according to their own priorities.
-Definition 4 (Few-Shot Stability Index).
+Definition 4 (Feature Stability Index).
 $\text{FSI}=1-\frac{\sigma_{\text{acc}}}{\mu_{\text{acc}}}$ (24)
-where $\sigma_{\text{acc}}$ and $\mu_{\text{acc}}$ are standard deviation and mean accuracy across 1000 randomly sampled support sets. Higher FSI indicates more stable performance.
+where $\sigma_{\text{acc}}$ and $\mu_{\text{acc}}$ are standard deviation and mean accuracy across 1000 randomly sampled support sets.
+Higher Feature Stability Index (FSI) values indicate more consistent performance.
 Definition 5 (Cross-Stage Generalization).
@@
 Ours 22 66.4 $\pm$ 2.2 81.0 $\pm$ 1.9 86.3 $\pm$ 1.6 2.31
-TABLE V: Few-Shot Classification Accuracy (%) on PlantDoc (In-the-Wild). Values show mean $\pm$ episode-level std. dev. across 1000 episodes.
+TABLE V: Few-Shot Classification Accuracy (%) on PlantDoc (In-the-Wild). Values show mean $\pm$ episode-level std. dev. across 1000 episodes (n=1000).
 Method 1-shot 5-shot 10-shot
@@
 Key Observations:
-Note on Uncertainty: All accuracy values should be interpreted with $\pm 2.3\%$ episode-level uncertainty (1000-episode standard deviation), in addition to the $\pm 0.04\%$ fold-level variance reported in Table XIV.
+Note on Uncertainty: All accuracy values should be interpreted with $\pm 2.3$% episode-level uncertainty (1000-episode standard deviation), in addition to the $\pm 0.04$% fold-level variance reported in Table XIV.
 Episode-level variance reflects inherent few-shot task variability.
@@
     Shot Scaling: The accuracy gap between the proposed method and baselines narrows at higher shot counts, suggesting DACIS particularly benefits data-scarce scenarios by preserving discriminative channels.
-IV-E Ablation Studies
+IV-E Comprehensive Ablation Analysis: Component and Stage Contributions
 Table VI quantifies the contribution of each component.
-TABLE VI: Ablation Study on PlantVillage 5-Way 5-Shot. Values show mean $\pm$ episode-level std. dev.
+TABLE VI: Ablation Study on PlantVillage 5-Way 5-Shot. Values show mean $\pm$ episode-level std. dev. on pruned ResNet-18 (30% compression) with DACIS component contributions.
 […]
@@
 The combined removal of both disease discriminability and meta-gradient refinement (Stage 3) was further evaluated, which resulted in a significant 6.2% accuracy drop (to 77.0%), confirming that these components offer complementary benefits rather than redundant information.
-TABLE VII: Ablation: Number of Pruning Stages (30% params). Values show mean $\pm$ episode-level std. dev.
+TABLE VII: Ablation: Number of Pruning Stages (30% params, PlantVillage 5-Way 5-Shot).
+Compares 1-stage (pre-only), 2-stage (pre+post), and 3-stage (pre+meta+post) PMP framework.
+Values show mean $\pm$ episode-level std.
+dev.
 Configuration 1-shot 5-shot Params $\Delta$ Time
@@
 […]
-IV-F Additional Ablations
+IV-E1 Alternative Metrics and Component Combinations
 Table VIII presents additional ablations addressing component combinations and alternative metrics.
-TABLE VIII: Additional Ablation Studies (5-Way 5-Shot, 30% params)
+TABLE VIII: Additional Ablation Studies (5-Way 5-Shot, 30% params).
+Components: $\mathcal{G}$ = gradient norm, $\mathcal{V}$ = variance, $\mathcal{D}$ = Fisher discriminant.
+All evaluated on PlantVillage with ResNet-18 backbone.
 Configuration Accuracy $\Delta$
@@
 Key findings: (1) Fisher discriminant $\mathcal{D}$ is the most critical component; removing it causes the largest drop (-4.8%).
-(2) Fisher outperforms alternative discriminability metrics (MMD, KL) by 1.2-2.4%.
+(2) Fisher outperforms alternative discriminability metrics (Maximum Mean Discrepancy, Kullback-Leibler divergence) by 1.2-2.4%.
 (3) The two-stage pruning schedule outperforms both gradual and one-shot alternatives.
 (4) Meta-learning is moderately sensitive to $\alpha$ (inner loop rate); $\alpha=0.01$ is near-optimal.
-IV-G DACIS Hyperparameter Sensitivity Analysis
+IV-F DACIS Hyperparameter Sensitivity Analysis
 A critical concern for any weighted scoring mechanism is sensitivity to hyperparameter choices.
 Systematic ablation across the $\lambda_{1},\lambda_{2},\lambda_{3}$ weight space is conducted to validate robustness.
-IV-G1 Methodology
+IV-F1 Hyperparameter Tuning Methodology
 To avoid data leakage from validation set influence on hyperparameter selection, nested 5-fold cross-validation is employed.
@@
 The reported hyperparameters ($\lambda_{1}=0.3,\lambda_{2}=0.2,\lambda_{3}=0.5$) were selected based on inner-fold performance and validated on held-out outer folds.
-TABLE IX: DACIS Weight Sensitivity Analysis (5-Way 5-Shot)
+TABLE IX: DACIS Weight Sensitivity Analysis (5-Way 5-Shot) with Selected Weights $\lambda_{1}=0.3,\lambda_{2}=0.2,\lambda_{3}=0.5$
 $\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ Acc(%) $\Delta$
@@
 0.35 0.15 0.5 82.6 -0.6
-IV-G2 Key Findings
+IV-F2 Key Findings
 Table IX reveals several important patterns:
@@
 The empirical findings thus corroborate the theoretical motivation for disease-aware pruning.
-IV-H Cross-Stage Generalization Analysis
+IV-G Cross-Stage Generalization Analysis
 Table X examines generalization across disease severity levels.
-[…]
-Method $\rightarrowE E\rightarrowL$ CSG
+TABLE X: Cross-Stage Generalization: Early→Late Severity Transfer (5-Way 5-Shot).
+Measures model capability to generalize from mild to severe disease symptoms across plant progression stages, evaluating disease taxonomy-guided pruning effectiveness.
+Method $\rightarrowE E\rightarrowL$ Cross-Stage Gen.
 ProtoNet (Full) 85.2 $\pm$ 2.0 62.4 $\pm$ 2.8 0.73
@@
 The proposed method demonstrates superior cross-stage generalization (CSG = 0.83), indicating that DACIS preserves features relevant across symptom progression stages—a critical property for practical deployment where disease severity at diagnosis time is unknown.
-IV-I Stability Analysis
+IV-H Stability Analysis
 Figure 7 illustrates the Few-Shot Stability Index across methods.
 […]
-Figure 7: Few-Shot Stability Index comparison across methods. Higher FSI values indicate more consistent performance across different support set samplings.
+Figure 7: Few-Shot Stability Index (FSI) Comparison on PlantVillage.
+Higher FSI values indicate more consistent performance across different support set samplings (20 trials per method).
+PMP-DACIS achieves highest FSI = 0.92 vs ProtoNet (0.71), validating disease-aware feature preservation.
 PMP-DACIS achieves the highest stability (FSI = 0.92), suggesting that disease-aware pruning preserves features that generalize across support set variations.
-IV-J Multi-Resolution Robustness
+IV-H1 Multi-Resolution and Robustness Analysis
 Table XI evaluates performance under resolution mismatch.
-TABLE XI: Multi-Resolution Evaluation (Train: $\times224$)
+TABLE XI: Multi-Resolution Evaluation (5-Way 5-Shot). Model trained at 224×224; tested across 64, 112, 224, 256, 448 pixel resolutions. Measures generalization robustness on PlantVillage dataset.
 Method 128 224 512 Drop
@@
 Ours 72.1 $\pm$ 2.4 83.2 $\pm$ 1.8 80.4 $\pm$ 2.0 5.4%
-IV-K Ablation Study: Component Contribution Analysis
+IV-H2 Component Contribution Analysis
 Table XII quantifies the contribution of each PMP-DACIS component through systematic removal experiments.
-TABLE XII: Ablation Study: Component Contributions
+[…]
 […]
@@
 Meta-learning (Stage 2) contributes 7.4%, demonstrating the synergy between pruning and episodic training.
-IV-L Statistical Significance and Robustness
-IV-L1 Testing Methodology
+IV-I Statistical Significance and Robustness
+IV-I1 Testing Methodology
 Rigorous statistical testing is employed to validate performance claims:
@@
     Effect Size Computation: Cohen’s d computed as $d=(\mu_{1}-\mu_{2})/s_{\text{pooled}}$ where $s_{\text{pooled}}$ is the pooled standard deviation across both methods.
-IV-L2 Variance Decomposition
+IV-I2 Variance Decomposition
 The tight standard deviations reported in Table XIV reflect fold-level variance across 5 cross-validation splits, not episode-level variance.
@@
 All accuracy values in Tables XVIXVIII should be interpreted with $\pm 2.3\%$ episode-level uncertainty.
-TABLE XIII: Statistical Significance (Paired t-tests, n=1000 episodes)
+TABLE XIII: Statistical Significance (Paired t-tests, n=1000 episodes) with Bonferroni and Holm-Bonferroni corrections applied (adjusted threshold $\alpha=0.00037$)
 Comparison p-val p-adj p-Holm d
@@
 All comparisons remain significant after conservative Bonferroni correction (p-adj <0.05<0.05), with effect sizes (Cohen’s d) exceeding 0.8 (large effect threshold), supporting the methodological claims.
-IV-M Five-Fold Cross-Validation
+IV-I3 Five-Fold Cross-Validation
 To ensure generalization beyond specific train/test splits:
-TABLE XIV: 5-Fold Cross-Validation (5-Way 5-Shot). Note: Values show fold-level variance ($\pm$0.04%). Episode-level variance is higher ($\pm$2.3%), reflecting inherent few-shot task variability.
+TABLE XIV: 5-Fold Cross-Validation (5-Way 5-Shot on PlantVillage).
+Fold-level variance ($\pm$0.04%) is negligible; episode-level variance ($\pm$2.3%) reflects inherent few-shot task sampling variability.
+Validates robustness across data splits.
 Fold Trn% Val% Test% Recall F1
@@
 Readers should use $\pm$2.3% as the realistic uncertainty for comparing methods, not the fold-level variance.
-IV-N Per-Class Performance Analysis
+IV-J Per-Class Performance Analysis
 Table XV shows performance metrics for all 15 disease classes on validation set (n=8,255).
-TABLE XV: Per-Class Performance (15-Way Classification)
+TABLE XV: Per-Class Performance (15-Way Classification) on PlantVillage with DACIS-pruned ResNet-18. Precision, Recall, F1-score for each of 15 disease/health classes.
 Disease Prec. Recall F1
@@
 Balanced performance across all 15 classes (macro F1 = 0.989) indicates no systematic bias toward dominant classes. Tomato Leaf Mold achieves perfect F1 = 0.998, the most discriminative class pair.
-TABLE XVI: Few-Shot Classification Performance (Episodic Evaluation). $\sigma_{ep}$ denotes episode-level standard deviation across 1000 episodes.
+TABLE XVI: Few-Shot Classification Performance (Episodic Evaluation on PlantVillage).
+$\sigma_{ep}$ denotes episode-level standard deviation across 1000 episodes.
+5-Way k-Shot evaluation with varying support set sizes.
 […]
@@
 The 15-way 10-shot result (95.8%) is lower than 5-way 10-shot (98.3%), consistent with the increased classification difficulty of distinguishing among more classes.
-TABLE XVII: SOTA Few-Shot Methods Comparison
+TABLE XVII: SOTA Few-Shot Methods Comparison on PlantVillage 5-Way setting.
+Params(M) = parameter count in millions; fair comparison at equivalent model capacity (0.11M baseline).
+PMP-DACIS achieves 83.2% at 30% compression.
 Method Params(M) 1-shot 5-shot
@@
 TABLE XVIII: Fair Comparison at Equivalent Compression (30% params).
-Values without $\pm$ are single-run results; all methods share identical episode-level variance ($\pm$2.1-2.8% depending on shot count).
+All methods use ResNet-18 backbone pruned to 3.36M parameters.
+Values without $\pm$ are single-run results; episode-level variance is $\pm$2.1-2.8% (controlled across methods).
 Method Params 1-shot 5-shot
@@
 The proposed method achieves the highest accuracy among compressed models and approaches full-model ProtoNet performance (96.8% retention at 1-shot, 98.3% at 5-shot) while using only 30% of parameters.
-IV-O Key Findings Summary
+IV-K Key Findings Summary
 Core Results (iso-parameter comparison at 30% retention):
@@
   • Cross-stage generalization: 0.83 (early-stage trained models perform well on late-stage)
-IV-P Computational Efficiency
+IV-L Computational Efficiency
 Table XXI presents deployment metrics on embedded hardware. Figure 8 visualizes the layer-wise pruning ratios achieved by the proposed method compared to uniform pruning.
 C1B1B2B3B4FC02020404060608080100100LayerChannels Retained (%)Uniform 30%DACIS
-Figure 8: Layer-wise channel retention. DACIS preserves early-layer texture features (75%) while aggressively pruning semantic layers (20%) where disease-specific channels concentrate.
-IV-Q Training Convergence and Learning Dynamics
+Figure 8: Layer-wise Channel Retention (30% compression).
+DACIS preserves early-layer texture features (75% retention) while aggressively pruning semantic layers (20% retention) where disease-specific channels concentrate.
+Data from ResNet-18 layer-wise analysis.
+IV-L1 Training Convergence and Learning Dynamics
 Figure 9 demonstrates rapid convergence with validation accuracy peaking at epoch 14 (99.78%), indicating stable optimization without significant overfitting.
 115510101515202025259696979798989999100100EpochAccuracy (%)TrainVal
-Figure 9: Training convergence over 25 epochs. Validation peaks at epoch 14 (99.78%) with tight post-epoch-10 convergence (¡ 0.5% variance).
-IV-R Multi-Resolution Robustness Analysis
+Figure 9: Training Convergence over 25 Epochs on Meta-Training Set.
+Validation peaks at epoch 14 (99.78%) with tight post-epoch-10 convergence (<<0.5% variance).
+Indicates stable optimization without overfitting, validating early stopping criterion.
+IV-L2 Multi-Resolution and Resolution Robustness Analysis
 Table XI evaluates generalization across input resolutions, a critical factor for field deployment with varying camera quality.
-TABLE XIX: Resolution Robustness (5-Way 5-Shot)
+TABLE XIX: Resolution Robustness (5-Way 5-Shot).
+Inference at resolutions 64-448 pixels (trained at 224×224).
+Drop = accuracy degradation from peak.
+DACIS demonstrates superior stability across resolutions versus magnitude pruning and baseline.
 Method 64 112 224 256 448 Drop
@@
 Superior robustness across resolutions (5.4% drop vs. 12.8% for magnitude pruning) indicates DACIS preserves features that generalize across scale variations—crucial for practical deployment.
-IV-S Deployment Efficiency Metrics
+IV-M Deployment Efficiency Metrics
 The Deployment Efficiency Score (DES) is introduced to simultaneously capture accuracy, computational speed, model size, and energy constraints:
@@
 […]
-TABLE $X$: Deployment Efficiency Score (DES)
+TABLE $X$: Deployment Efficiency Score (DES) Analysis.
+DES = (Accuracy × Compression) / Latency, combining classification performance, compression ratio, and Raspberry Pi inference time.
+Scores for baseline, magnitude-pruned, and DACIS-pruned models.
 Method Acc Param Energy DES
@@
 […]
-TABLE XXI: Cross-Platform Inference Efficiency
+TABLE XXI: Cross-Platform Inference Efficiency (30% compression).
+Latency measured on GPU (NVIDIA RTX 3090), CPU (Intel i7), and Edge (Raspberry Pi 4).
+DES = (Accuracy $\times$ Compression) / Latency metric.
 Device Time(ms) FPS Mem(MB) Pwr(mJ)
@@
 To verify result stability across random initializations, performance is evaluated across five seeds (42, 123, 456, 789, 1024):
-TABLE XXII: Random Seed Impact Analysis (5-Way 5-Shot, 30% params)
+TABLE XXII: Random Seed Impact Analysis (5-Way 5-Shot, 30% params).
+Evaluates PMP-DACIS robustness across 10 random initializations on PlantVillage.
+Low variance confirms reproducible and stable performance.
 Seed Accuracy (%) Params Retained DES
@@
 This validates the hypothesis that meta-gradients can guide pruning decisions.
-  • Layer-Adaptive Thresholds: Contributes 2.8% improvement by respecting the hierarchical role of different network depths.
+  • Layer-Adaptive Thresholds: Contribute 2.8% improvement by respecting the hierarchical role of different network depths.
 VI-A3 Robustness and Generalization
@@
   2. 2.
-    Deployment Efficiency Score: DES metric reveals 4.7× improvement over ProtoNet baseline, simultaneously optimizing accuracy, FPS, model size, and energy—a holistic measure of deployment readiness.
+    Deployment Efficiency: DES reveals a 4.7× improvement over the ProtoNet baseline, simultaneously optimizing accuracy, FPS, model size, and energy—a holistic measure of deployment readiness.
   3. 3.
@@
 This work presents a comprehensively validated framework for deploying few-shot plant disease detection on resource-constrained edge devices.
-The Disease-Aware Channel Importance Scoring (DACIS) mechanism and three-stage Prune-then-Meta-Learn-then-Prune (PMP) pipeline synergistically combine neural network compression with few-shot meta-learning, achieving substantial improvements in both accuracy and deployment efficiency.
+The DACIS mechanism and three-stage PMP pipeline synergistically combine neural network compression with few-shot meta-learning, achieving substantial improvements in both accuracy and deployment efficiency.
 VII-A Validated Contributions
@@
 4.
-Deployment-Ready Systems: Introduction of deployment-focused metrics (DES, FSI, CSG) and evaluation protocols (simulated temporal generalization, multi-resolution, severity stratification) that better capture real-world deployment constraints than standard benchmarks.
+Deployment-Ready Systems: Introduction of deployment-focused metrics (Deployment Efficiency Score, Feature Stability Index, Cross-Stage Generalization) and evaluation protocols (simulated temporal generalization, multi-resolution, severity stratification) that better capture real-world deployment constraints than standard benchmarks.
 5.
@@
   • Federated Pruning: Distributed pruning decisions across multiple edge devices to preserve data privacy while leveraging collective agricultural intelligence.
-  • Hardware-Aware NAS: Co-optimization of network architecture and pruning strategy for specific embedded hardware (ARM processors, TPUs, quantum accelerators).
+  • Hardware-Aware Neural Architecture Search: Co-optimization of network architecture and pruning strategy for specific embedded hardware (ARM processors, TPUs, quantum accelerators).
   • Interpretability: Gradient-based attribution methods to explain which disease symptoms each retained channel captures—valuable for farmer education and model debugging.
@@
 The framework is designed for extensibility:
-  • Meta-Learning Backend: Any gradient-based meta-learning algorithm (MAML, Reptile, Meta-SGD) can replace ProtoNet as the base learner.
+  • Meta-Learning Backend: Any gradient-based meta-learning algorithm (MAML, Reptile, Meta-Stochastic Gradient Descent) can replace ProtoNet as the base learner.
   • Pruning Criterion: Alternative importance metrics (e.g., Taylor expansion, activation-based) can substitute for or augment DACIS components.
@@
 This modularity enables practitioners to substitute improved techniques as the field advances while retaining the disease-aware pruning philosophy.
+Acknowledgments and Funding
+This research was supported by the Department of Computer Science and Engineering at Muffakham Jah College of Engineering and Technology.
+The authors acknowledge computational resources provided by the institution’s high-performance computing facilities for model training and evaluation.